### v2

# WallFacer: Guiding Transformer Model Training Out of the Long-Context Dark Forest with N-body Problem

Abstract.
In recent years, Transformer-based Large Language Models (LLMs) have garnered significant attention due to their exceptional performance across a variety of tasks. However, training these models on long sequences presents a substantial challenge in terms of efficiency and scalability. Current methods are constrained either by the number of attention heads, limiting scalability, or by excessive communication overheads. In this paper, we propose an insight that Attention Computation can be considered as a special case of n-body problem with direct interactions. Based on this concept, this paper introduces WallFacer, an efficient long-sequence training system with a novel multi-dimensional ring sequence parallelism, fostering an efficient communication paradigm and extra tuning space for communication arrangement. Through comprehensive experiments under diverse environments and model settings, we demonstrate that WallFacer significantly surpasses state-of-the-art method that supports near-infinite sequence length, achieving performance improvements of up to 77.12%.
distributed deep learning, sequence parallelism, n-body simulation, large scale training, high performance computing
copyright: noneconference: arxiv; Mar, 2025; ROTTERDAM
1. Introduction
Over the past decade, deep learning has made remarkable strides in diverse fields, including computer vision (CV) and natural language processing (NLP). Transformer-based (Vaswani et al., 2023) models, leveraging the groundbreaking attention mechanism introduced in 2018, have secured a dominant position due to their superior feature-capturing capabilities. As the technology has evolved, the ability to efficiently process long sequences has emerged as a pivotal challenge, capturing the attention of researchers and practitioners alike.
Figure 1. Comparison of the theoretical peer-to-peer communication amount for Ring Attention and different configurations of WallFacer while training a llama-7B model(Touvron et al., 2023) on 64 GPUs with Adam optimizer(Kingma and Ba, 2017) and a batch size of 4.
This focus on long-sequence training and inference is underscored by its critical importance in a variety of downstream tasks. For instance, in text summarization, the ability to handle extensive sequences is vital, as the content to be summarized can range from lengthy chapters to entire books (Koh et al., 2022; Beltagy et al., 2020). Similarly, chat-based applications, such as ChatGPT (Achiam et al., 2023), require the capacity to process extensive dialogue histories to ensure conversational consistency. There are also applications in other fields like video generation(Brooks et al., 2024; Peebles and Xie, 2023) and protein structure prediction(Jumper et al., 2021; Cheng et al., 2023). In summary, the demand for efficient long-sequence training and inference mechanisms spans a broad spectrum of applications, highlighting a critical area of need within the community.
The long context in the above scenarios has introduced several challenges for model training and inference: 1) Efficiency and Adaptability. The challenge of efficiency is predominantly determined by computation and communication. How to maintain high efficiency in various environments and settings has drawn great attention from academia and industry. 2) Memory. As the size of the activation during attention is quadratically related to the length of the sequence, memory consumption increases rapidly with the growth of the sequence. 3) Scalability. Large Language Models (LLMs) necessitate the use of thousands of GPUs for training, even with datasets of regular lengths. For long sequences, ensuring an acceptable scaling speedup rate is even more critical to reduce time and economic costs.
Traditional parallelisms such as Data Parallelism(Hillis and Steele Jr, 1986; Shoeybi et al., 2019; Li et al., 2020; You et al., 2018), Tensor Parallelism(Shoeybi et al., 2019; Wang et al., 2022; Xu et al., 2021), and Pipeline Parallelism(Huang et al., 2018; Fan et al., 2020; Li and Hoefler, 2021; Liu et al., 2023a) have not been able to address the large memory requirement of extremely long sequences. To break through this obstacle, Sequence Parallelism has been introduced, splitting the input on the sequence length dimension. Mainstream Sequence Parallelism schemes can generally be classified into two categories: those based on all-to-all communication, and those based on ring peer-to-peer communication. Methods like DeepSpeed Ulysses(Jacobs et al., 2023), which are based on all-to-all communication, offer efficiency but require the splitting of attention heads. Consequently, these methods are limited in scalability. On the other hand, peer-to-peer communication methods(Liu et al., 2023b; Li et al., 2024), such as Ring Attention(Liu et al., 2023b), do allow for infinite context lengths; however, they necessitate the transmission of complete keys and values across all GPUs, leading to significantly high communication loads. In summary, there remains a deficiency in communication-efficient methods that are capable of supporting infinite context lengths.
In fact, we can draw valuable insights from other fields that have extensive experience in handling long sequences. In the field of n-body simulations, researchers have been dealing with vast numbers of particles, typically far exceeding the number of tokens addressed in current Transformer models. Through careful observation, we have identified that the core component of Transformers, the attention mechanism, closely resembles a special case of the n-body problem. Specifically, n-body problems with direct interactions focus on calculating the resultant force exerted by all other particles in the system on each individual particle. Similarly, in self-attention, the objective is to compute the attention score that each token in the sequence assigns to every other token. This similarity suggests that we can leverage methodologies from n-body problem research—refined over decades—to inform and enhance our approaches to tackling the relatively new challenges of long-context Transformer training.
Inspired by the methodology of n-body communication optimization(Driscoll et al., 2013) and tailored to the Transformer architecture, we introduce WallFacer, a near-infinite-context Transformer training system with multi-ring sequence parallelism that incorporates an additional parallel dimension to the existing ring-style communication, fostering an efficient communication paradigm and extra tuning space for communication arrangement. With very little extra cost of memory, WallFacer parallelism significantly reduces the peer-to-peer communication amount, as is shown in figure 1.
In summary, our paper presents these contributions:
  • We conceptualize Attention computation as a novel instance of the traditional n-body problem, providing fresh insights into optimizing and parallelizing Attention computation.
  • We introduce a near-infinite-context training system for Transformer models, featuring a groundbreaking multi-ring sequence parallelism scheme. This scheme adds an additional dimension of parallelism and significantly reduces peer-to-peer communication, all while maintaining a minimal memory footprint.
  • We offer a straightforward method that allows users to select the most suitable parallelism scheme based on their specific needs, maximizing the utility of the available tuning space within our communication framework.
  • We perform experiments on mainstream Transformer models, conducting performance and scaling tests across various computing clusters. Preliminary results indicate that our WallFacer system outperforms Ring Attention by up to 77.12%, showcasing its efficacy and scalability.
2. Background
2.1. Transformer and Long Sequence Training
The key mechanism behind Transformer-based models is Attention(Vaswani et al., 2023), which captures the text feature by calculating the attention score between every two single tokens. A standard attention function is given as:
(1) $Attention(Q,K,V)=softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V$
where Q, K, and V are the Query, Value, and Key of input x with shape (B, N, H) in the case of self-attention, and $d_{k}$ represents the head dimension in commonly used multi-head attention. For convenience, we use the symbolic representations in Table 1. We can observe that in the original attention function, we need to store the intermediate value $QK^{T}$ of size $N^{2}$, which means that the memory capacity requirement and computation amount both grow in quadratic proportion with the sequence length. To solve the memory IO and capacity bottleneck, methods like flash-attention(Dao et al., 2022) have been proposed. Flash-attention is designed upon the idea of softmax decomposition (Kitaev et al., 2020; Milakov and Gimelshein, 2018; Rabe and Staats, 2022), enabling online softmax computation by storing some extra statistics. Flash-attention is thus capable of segmenting the Q, K, and V matrices into blocks, allowing the entire computation to be performed within a single kernel for each block sequentially. This strategy effectively circumvents redundant read and write operations to the High Bandwidth Memory (HBM). However, when the sequence length reaches millions, it becomes necessary to distribute the sequence across multiple GPUs. This distribution helps to reduce both the memory and computation demands on any single device. This strategy is also known as Sequence Parallelism. We will explore these methods in further detail in the following section.
Table 1. Meanings of the symbols that are used in this paper
P The number of GPUs
C The parallel size of WallFacer (attention parallel size)
H The hidden dimension size of the Transformer blocks
N The total number of tokens within the whole sequence
B The training batch size
W The communication bandwidth between GPUs
L The communication latency between GPUs
2.2. Sequence Parallelism
Sequence Parallelism was first conceptualized in a study published in 2021 (Li et al., 2021), which introduced the concept as a strategy to distribute computational sequences across multiple processing steps. In this paper, we adopt a broad definition of Sequence Parallelism to include any approach that partitions a sequence during specific computational phases. Presently, Sequence Parallelism can be divided into two main categories: attention-head-sharding-based and peer-to-peer-communication-based. The former involves distributing the attention heads of multi-head attention across multiple GPUs, whereas the latter resembles a distributed version of flash-attention, relying on peer-to-peer communication to transfer keys, values, and intermediate statistics. We will now explore the advantages and disadvantages of these methods in detail.
2.2.1. Attention-head-Sharding-Based
Here we introduce the two representative methods, DeepSpeed Ulysses and Megatron Sequence Parallelism.
DeepSpeed Ulysses. DeepSpeed Ulysses(Jacobs et al., 2023) presents a strategy for training with long sequences by leveraging all-to-all collective communication, but is largely limited in its scalability. Functions outside the attention layer are partitioned along the sequence dimension using conventional methods. Within the attention block, Ulysses transitions from sequence parallelism to a method akin to tensor parallelism. It divides the query, key, and value matrices across the attention heads, thereby preserving the original attention computation structure. This is facilitated by two sets of all-to-all communication that switch between sequence splitting and attention head spliting.
The principal merit of Ulysses lies in its simplicity and ease of implementation. Nonetheless, its reliance on the number of attention heads for partitioning activations introduces the following limitations. 1) Ulysses is limited in scalability, as it can only be expanded to as many GPUs as there are attention heads, which in turn restricts the maximum sequence length that can be processed. This limitation becomes particularly problematic when employing techniques like grouped-query attention (GQA) (Ainslie et al., 2023) or multi-query attention (MQA) (Shazeer, 2019), which further reduce the number of available attention heads for Keys and Values. 2) Ulysses encounters load-balancing issues when the number of attention heads is not evenly divisible by the number of GPUs, complicating its deployment across diverse hardware configurations. 3) The need for head-splitting makes Ulysses hard to combine with Tensor Parallelism.
Megatron Sequence Parallelism. Megatron Sequence Parallelism(Korthikanti et al., 2022) shares its name with an earlier study(Li et al., 2021). Megatron Sequence Parallelism focuses on minimizing memory usage and reducing the necessity for activation recomputation. The training approach is based on Tensor Parallelism (TP). While TP is employed in the Linear and Self-Attention blocks, operations such as the LayerNorm and Dropout functions are not distributed, leading to replications within the tensor parallel group. This requires an additional all-gather operation and a further reduce-scatter operation during the forward propagation to alternate between Tensor Parallelism and Sequence Parallelism. Megatron Sequence Parallelism is not designed for training with exceedingly long sequences, as the Self-Attention and MLP layers continue to rely on Tensor Parallelism and must process the full sequence length. Like DeepSpeed Ulysses, it is also limited by the number of attention heads.
2.2.2. Ring-peer-to-peer-communication-based
The primary method in peer-to-peer-communication-based strategies is Ring Attention(Liu et al., 2023b).
Ring Attention. Introduced in 2023, Ring Attention(Liu et al., 2023b) innovatively partitions the sequence dimension and utilizes a ring-style peer-to-peer (P2P) communication pattern to transfer Keys and Values across all GPUs. Each GPU receives the key and value matrices from the preceding rank, updates the local attention score, and then forwards them to the next rank, as is shown in Figure 2. This method employs an online-softmax and updates attention scores incrementally, allowing the computation of attention scores without retaining the full sequence length. Thus, it potentially supports infinite context, provided sufficient computing resources are available. However, the requirement for P steps of P2P communication renders this approach less efficient in environments with high-latency communication. A more detailed discussion on this topic is presented in the subsequent section.
Figure 2. An example of Ring Attention Computation on 16 GPUs in two nodes. 16 rounds of ring P2P communication are needed to reach the final results.
2.3. n-body Problem
Before diving into the model formulation, it is essential to have a brief overview of the n-body simulation. N-body simulations(Hockney and Eastwood, 1989) play a crucial role in the study and modeling of dynamics within multi-body systems. Typically, the N in n-body represents the number of entities, often reaching up to $10^{12}$, far exceeding the processing capability of a single processor. Consequently, n-body simulations are predominantly conducted on supercomputing clusters. Specifically, n-body problems involving direct interaction refer to scenarios where only the direct forces between particles, such as gravitational forces, are considered. The design of n-body simulations encounters several significant challenges similar to those discussed in previous sections: 1) Efficiency: Continuous efforts are directed towards developing more efficient frameworks and algorithms to enhance both the computational and communicational aspects of n-body simulations. 2) Memory: The number of particles in n-body problems typically surpasses those in natural language processing by orders of magnitude, presenting substantial challenges in managing memory on processors. 3) Scalability: As n-body simulations are often executed on a large amount of computing nodes, scaling the computations to more processors is notably challenging. This backdrop of challenges in n-body simulations leads us to ponder whether the decades of accumulated knowledge in this field could be leveraged to inform and guide the training of Transformer models.
3. Formulation: Transformer as N-body
Through observation, we have discerned that attention computation can be viewed as a new instance of the n-body problem with direct interactions. Moreover, the forward and backward propagation processes in neural networks can be interpreted as distinct scenarios within this n-body framework. This insight suggests that the methodologies developed for n-body problems could also be effectively applied to the long-sequence training of Transformer models.
3.1. Modeling of Attention as n-body
The general n-body problem with direct interaction can be described as:
(2) $B(x_{i})=\Sigma_{j=0}^{N}g(m_{0}(x_{j}),m_{1}(x_{i}))$
Taking the cosmology gravity problem as an example, $B(x_{i})$ represents the total gravitational force experienced by an astronomical object x within the system, $m_{0}(x)$ and $m_{1}(x)$ both denote the mass of $x$ in this case, $g$ refers to the gravitational formula between two objects, and $\Sigma$ denotes the formula for the resultant force composition. Similarly, attention computation in neural networks can be expressed in an analogous format. Moreover, this formulation allows for distinct interpretations of the processes during forward and backward propagation: Forward Propagation. From the nature of matrix multiplication and formula 1, we can tell that the forward attention computation is independent for each $Query_{i}$, and each $Query_{i}$ needs to calculate score for the $Key$s and $Value$s of every token. In this way, we can rewrite formula 1 into formula 2 by following the Flash-Attention(Dao et al., 2022) style and having:
(3) $\begin{split}m_{0}(x)&=W_{Q}x+B_{Q}=Q_{x}\\ m_{1}(x)&=W_{K}x+B_{K},\ W_{V}x+B_{V}=K_{x},V_{x}\\ g(m_{0}(x_{i}),m_{1}(x_{j}))&=softmax(\frac{Q_{xi}K^{T}_{xj}}{\sqrt{d_{k}}})V_{xj}\\ B(x_{i})&=\Sigma_{j=0}^{N}\ g(m_{0}(x_{i}),m_{1}(x_{j}))*e^{lse_{j}-lse_{new}}\end{split}$
where the Ws and Bs refer to the weights and biases in the Linear functions, while lse is the LogSumExp (Nielsen and Sun, 2016) result calculated during $g$.
Backward Propagation The primary distinction between forward and backward attention lies in the computational process during backward propagation. Specifically, to calculate the gradients for Keys and Values, it is necessary to iterate through all the Queries, and vice versa. This means that each token must interact with all three properties (Queries, Keys, Values) of every other token to compute its final gradients. We can analogize this to an n-body problem, where the interaction of forces is asymmetrical in two directions. Consequently, this requires conducting two rounds of the process as described in formula 2. In the first round, $x_{i}$ refers to $K_{i}$ and $V_{i}$ while in the second round, it refers to $Q_{i}$.
In summary, we can easily interpret both the forward and backward attention calculation into a simple form of the n-body problem.
Figure 3. An overview of the WallFacer Training System
3.2. Methodology Transfer
It is a recognized practice to harness the knowledge from traditional systems and high-performance computing to guide the design of machine learning systems. Now that we have modeled the attention calculation in Transformers as an n-body problem, we are poised to transfer a wealth of optimization methodologies from n-body simulation to Transformer training. This methodology transfer is deemed rational, as evidenced by existing techniques shared between these fields. For instance:
Particle Decomposition: Commonly used to manage memory and computation constraints, this strategy involves distributing particles across multiple processors and integrating the results using the law of force composition. This method is widely implemented in models that employ online softmax, such as Flash-Attention(Dao et al., 2022).
Hierarchical Methods: Hierarchical structures, such as the Multilevel Fast Multipole Method(Song and Chew, 1995; Lu and Chew, 1994) or tree-based processes like the Barnes–Hut simulation(Barnes and Hut, 1986), are used to approximate n-body calculations. These methods reduce the complexity from $O(n^{2})$ to $O(n\log n)$ or $O(n)$. Such approaches have been effectively applied in various fields, including computer vision, as seen in FASTERVIT(Hatamizadeh et al., 2024), etc., and in natural language processing tasks, as demonstrated in (Liu and Lapata, 2019) and so on.
Cutoff: Interactions in many systems decay with distance, making it practical to set a cutoff distance and ignore interactions beyond this threshold. This principle is similarly utilized in approximate attention algorithms like Longformer(Beltagy et al., 2020), which employs a sliding window to calculate an approximate attention score for nearby tokens while excluding those outside the window. These examples validate the feasibility of guiding Transformer system design using n-body simulation techniques. Furthermore, we have identified additional aspects of n-body simulation that could inspire innovative approaches in training long-sequence Transformer models.
In addition to the aforementioned techniques, we have explored a series of multi-dimensional algorithms used but not limited in n-body problems(Solomonik and Demmel, 2011; Aggarwal et al., 1990; Dekel et al., 1981; Agarwal et al., 1995; Gupta and Sadayappan, 1994) that optimize the communication scheme among processors while significantly reducing the total communication load, albeit at the expense of increased memory usage. The core idea involves leveraging an additional parallel dimension to partition the complete communication and computation tasks, distributing them across various devices. Inspired by these developments, we have conceived the concept of multi-ring sequence parallelism for our training system, which will be discussed in detail in the following section.
4. WallFacer Training System
4.1. Overview
Built upon the inspiration from n-body simulation, the WallFacer training system is comprised of five main components. At the core of the system is WallFacer Attention, which leverages multiple ring-style P2P (peer-to-peer) communication strategies to enhance the efficiency of distributed attention computation. The Dataloader is designed to organize tokens within each sub-sequence according to their mask (causal or full) and distribute them across different GPUs. The Communication Configuration generator plays a critical role in initially assigning Keys and Values to their corresponding ranks. Meanwhile, the Communication Topology Scheduler outlines the placement of parallelism across various computing nodes. Finally, the WallFacer Runtime incorporates additional supporting techniques for the training process, such as gradient checkpointing. We will now explore the details of these components in depth.
4.2. WallFacer Attention
Figure 4. An example of a Transformer block with WallFacer Attention on one team of four devices out of 64 GPUs.
As discussed in the previous section, a major limitation of Ring Attention is the extensive amount of peer-to-peer (P2P) communication required, which becomes problematic in environments with weak connections between computing nodes. Drawing inspiration from the multi-dimensional parallelism used in n-body simulations, we enhance the ring sequence parallelism by introducing an additional dimension. This is achieved by duplicating the Queries, Keys, and Values, thus dividing the communication tasks and distributing them within each team. This segmentation markedly improves computational efficiency and scalability.
We will now delve into the specifics of the WallFacer Parallelism and provide a theoretical analysis demonstrating its advantages over Ring Attention.
4.2.1. Training Process of WallFacer Parallelism
In the WallFacer system, GPUs are grouped into Teams to coordinate computation and communication tasks more effectively, unlike the isolated operations in traditional settings. The Queries, Keys, and Values are gathered within the team, so each member of the team holds the same activation of the whole team before attention. WallFacer introduces an additional parameter, C, which determines the replication factor of the input and, consequently, the number of GPUs within each team. The range of C is from 1 to $\sqrt{P}$. When C equals one, the algorithm falls back to Ring Attention. When C equals $\sqrt{P}$, the algorithm becomes a completely collective-communication-based one. When $1<C<\sqrt{P}$, it becomes a structure with multiple rings looping concurrently. There are three main distinctions between the operational processes of WallFacer and Ring Attention: QKV_matmul & all-gather, attention iteration, and reduce-scatter. Throughout the attention process, asynchronous communication is employed alongside the early launch of communication kernels to maximize the overlap of computation and communication tasks. The complete forward process of a WallFacer Transformer block is outlined in Algorithm 1.
Algorithm 1 WallFacer Attention Block (Forward)
0:  Input sequence x, Linear Function query, key, and value, attention parallelism size c, global rank r, global size gs, team process group pg
[…]
2:  compute the initial rank to send $\textbf{r}_{send}=$$get\_init\_send(\textbf{r})$
3:  compute the initial rank to receive from $\textbf{r}_{recv}=$$get\_init\_recv(\textbf{r})$
[…]
6:  initialize attention score O, extra statistics lse to zero.
7:  for $1\leq i\leq world\_size/c^{2}$ do
8:     wait for $\textbf{req}_{send}$ and $\textbf{req}_{recv}$
9:     $\textbf{k}_{current}=\textbf{k}_{next}$, $\textbf{v}_{current}=\textbf{v}_{next}$
10:     launch $\textbf{req}_{send}$ to send $\textbf{k}_{current}$ and $\textbf{v}_{current}$ to $\textbf{r}_{next}$, launch $\textbf{req}_{recv}$ to receive $\textbf{k}_{next}$ and $\textbf{v}_{next}$ from $\textbf{r}_{last}$
[…]
12:  end for
[…]
14:  return $\textbf{O}_{final}$
Forward Propagation. In Figure 4, we have an example of one team of four GPUs out of all the 64 GPUs performing WallFacer-style attention. Each training iteration begins with the dataloader splitting the entire input sequence of length N into N/P sub-sequences, which are then loaded onto each GPU. As previously mentioned, the next step involves computing the Queries, Keys, and Values. These are computed separately via matrix multiplication, followed immediately by the launch of the all-gather kernel, which gathers the above QKVs within the team, allowing for the overlap of up to two-thirds of the communication with computation.
Once this phase is complete, each GPU within the team possesses the same Q, K, and Vs, each of a length of $\frac{CN}{P}$. To distribute the communication and computation tasks among the team members, we divide the original workload based on four specific ranks assigned to each GPU by the Communication Configuration Generator. These ranks determine each GPU’s partners and position within the P2P ring, details of which will be discussed in Section 4.3.
Following the setup, the Keys and Values are dispatched to their designated locations within the cluster to establish the initial sub-ring, setting the stage for the multi-ring iteration phase of WallFacer attention. Given that each sub-sequence is $\frac{CN}{P}$ long and each GPU is tasked with computing the attention score for $\frac{1}{C}$ of the whole sequence, it results in $\frac{N/C}{CN/P}=P/C^{2}$ rounds of communication. This implies that there are $P/C^{2}$ GPUs in one ring.
The iteration process, conducted in a Flash-Attention style, involves storing the log-sum-exp (lse) and intermediate output O, which are updated step by step. Queries are retained locally, while Keys and Values circulate through the ring via P2P communication. After completing the iterations, each team member accumulates the attention scores for the entire team’s sub-sequence of Queries with 1/C of the Keys and Values from the full sequence.
A simple reduce-scatter operation is then employed to amalgamate the intermediate results and distribute them among the team members. Each GPU ultimately contains the final attention score for its portion of the sequence over the entire sequence. The output of this forward attention block is finalized after a standard LayerNorm and FeedForward layer process.
Backward Propagation. The major distinction between backward and forward propagation, as outlined in section 3, is the inability to calculate Queries independently during the backward phase. Unlike forward propagation, the backward phase requires the complete set of Keys and Values to calculate the gradient for Queries, and vice versa. To manage this, we have structured the gradient calculation into two loops: the Key & Value outer loop and the Query inner loop.
In the outer loop, gradients for Keys and Values are tracked and maintained fixed on the corresponding GPUs within the sub-rings; these gradients do not transfer between GPUs. The inner loop, however, handles the gradients for Queries, which start initialized as zero and are circulated along the sub-rings together with the Queries themselves. During each iteration, the approach mirrors the backward computation method used in Flash Attention, where the updated gradient of the current Query shard is passed to the next GPU in the ring, while the gradients for Keys and Values are retained for subsequent Query shards.
Initial communication ranks for sending and receiving are still configured by the Communication Configuration Generator. Additionally, there is an extra P2P communication required at the end of the process to send the Query and its gradient back to their original location following the loop.
4.2.2. Theoretical Analysis
In this section, we will discuss two major questions: how much communication reduction can WallFacer bring compared with Ring Attention, and does this come with a high extra memory cost? During the analysis, we will employ a case study using the WallFacer system with an attention parallel size of C = 4 on a llama-30B model, which consists of 64 layers. For this model, referred to as model M, the batch size = B is set to 1, the sequence length = N to 65,536, the hidden dimension = H to 6,656, and the number of GPUs = P to 64. Additionally, the computation will utilize bfloat16 precision.
Communication Analysis. Let’s analyze the communication overhead within one forward Transformer block on a single GPU. For Ring Attention, the communication is primarily due to the ring P2P loop. Each iteration’s communication overhead can be calculated as follows:
(4) $\frac{2BNH}{PW}+L$
and as the total number of iterations done is $P$, the total communication overhead will be:
(5) $\frac{2BNH}{W}+PL$
and this overhead can be partially overlapped with the attention computation.
For WallFacer, the communication overhead comes from collective and P2P both. The collective overhead for all-gather and reduce-scatter is:
(6) $\frac{4BNH(C-1)}{PW}$
while the P2P communication can be similarly computed as:
(7) $\frac{P}{C^{2}}(\frac{2CBNH}{PW}+L)=\frac{2BNH}{CW}+\frac{P}{C^{2}}L$
The advantages of WallFacer over Ring Attention during the ring-P2P phase are evident in three main aspects: 1) Reduced Communication and Latency: Ring Attention requires C times more communication than WallFacer, significantly increasing the bandwidth requirement across the entire cluster. For the llama 30B model M, the total communication volume of ring P2P communication and collective communication volume for Ring Attention and WallFacer can be computed as 1.625 GB and 0.152 GB(collective) + 0.406GB (P2P) = 0.558GB. Furthermore, while Ring Attention necessitates $P$ iterations per attention block, WallFacer only requires $\frac{P}{C^{2}}$, reducing the latency overhead by $C^{2}$. 2) Localized Communication: In scenarios like those depicted in Figure 3, WallFacer’s ring P2P communication can be confined within the same computing node, where bandwidth is typically much higher than between computing nodes. Conversely, Ring Attention demands inter-node communication during every iteration, which can be less efficient. 3) Enhanced Overlap of Communication and Computation: During each iteration, the communication volume of WallFacer is $C$ times higher than that of Ring Attention, while the computational volume during attention is approximately $C^{2}$ times greater. This higher computation-to-communication ratio makes it easier for WallFacer to overlap P2P communication with computation, enhancing overall efficiency.
Additionally, the collective communication in WallFacer is minimal due to: 1) Efficient Overlapping: The all-gather communication overlaps significantly with the QKV matrix multiplication, and the reduce-scatter communication partially coincides with the final attention score update. 2) Scale Considerations: Compared to the P2P communication column, there is a P in the denominator, which is substantial during large-scale training. This implies that collective communication constitutes a very small portion of the total communication volume.
Memory Analysis. In this section, we estimate the theoretical peak memory requirements necessary to store the model weights, activations, and optimizer states. Our implementation utilizes the Adam Optimizer (Kingma and Ba, 2017), bfloat16 precision, and Zero-2 optimization (Rajbhandari et al., 2020). Since the WallFacer architecture does not alter the model weights or the optimizer, we can assume that the memory costs associated with the model and the optimizer remain constant across both methods. We name the memory cost for the model and optimizer as $M_{m+o}$. As for the activation, we refer to the size of one single activation of a sub-sequence on one GPU as
(8) $A=\frac{B\times N\times H}{P}$
As we use the checkpointing scheme from (Li et al., 2024), a model of $Y$ layers needs to save $Y+1$ activations as checkpoints. Now we calculate the approximate peak memory after Q, K, and V are already calculated and before the attention computation at the last layer of the whole model. For Ring Attention and WallFacer, the peak memories are:
(9) $PM_{Ring}=M_{m+o}+(L+1)A+3A=M_{m+o}+(L+4)A$
(10) $PM_{Wall}=M_{m+o}+(L+1)A+3CA=M_{m+o}+(L+3C+1)A$
, where C is the WallFacer attention dimension. And for the example model M, the peak memory would be $M_{m+o}+68A$ and $M_{m+o}+77A$, and the extra memory cost compared with Ring Attention is less than 13.2%, while the P2P communication volume is reduced by about 75%. In a word, the extra memory cost is acceptable as a tradeoff for the communication reduction.
4.3. Communication Configuration Generator
The Communication Topology Scheduler, as introduced in the previous section, plays a crucial role in the initial setup of Queries, Keys, and Values on GPUs and determining each GPU’s position within the sub-rings.
Figure 5. An example of ring initialization process of 8GPUs and 4 sub-rings in the Communication Configuration Generator
Initially, for the setup stage, it is essential to establish the sub-rings by rearranging the activation positions. Specifically, during forward propagation, the Queries do not require rearrangement; however, the Keys and Values must be transmitted to their corresponding positions in the ring prior to commencing the loop. As illustrated in Figure 5, this initialization ensures that each team member holds a different shard of Keys and Values. Moreover, it guarantees that no two teams within the same ring possess identical Keys and Values.
Furthermore, Team 0 and Team 1 form what is termed a team group, where ring P2P communication occurs exclusively within this group, similarly for Team 2 and Team 3. A team group is defined as a collection of teams that participate in the same sub-rings. The number of teams within a team group can be calculated as $Team\ size=\frac{P}{C^{2}}$. We give the details of the determination of initial sending targets in the case of placing teams within same computing nodes in Algorithm 2, and the receiving source can also be calculated similarly.
Algorithm 2 get_init_send()
0:  inter-team rank $\textbf{r}_{t}$, intra-team rank $\textbf{r}_{a}$, inter-team dimension $\textbf{d}_{t}$, intra-team dimension $\textbf{d}_{a}$
1:  team group size = $\textbf{d}_{t}$ / $\textbf{d}_{a}$
2:  target team group rank = $\textbf{r}_{a}$
3:  target team = target team group rank * team group size + $\textbf{r}_{t}$ // $\textbf{d}_{a}$
4:  target device intra-team rank = $\textbf{r}_{t}$ $\textbf{d}_{a}$
5:  target global rank = target team * $\textbf{d}_{a}$ + target device intra-team rank
6:  return target global rank
After the initialization of activations, we can set up the rings by providing the GPUs their last and next GPU within their rings, as is described in Algorithm 3
Algorithm 3 get_P2P_config()
0:  inter-team rank $\textbf{r}_{t}$, intra-team rank $\textbf{r}_{a}$, inter-team dimension $\textbf{d}_{t}$, intra-team dimension $\textbf{d}_{a}$
1:  team group size = $\textbf{d}_{t}$ / $\textbf{d}_{a}$
2:  self team group rank = $\frac{r_{t}}{team\ group\ size}$
3:  next team in group = $(r_{t}+1)\%$ team group size $+$ team group size $\times$×self team group rank
4:  last team in group = $(r_{t}-1)\%$ team group size $+$ team group size $\times$×self team group rank
5:  next device global rank = $r_{a}+$ next team in group ×$\times d_{a}$
6:  last device global rank = $r_{a}+$ last team in group ×$\times d_{a}$
7:  return next device global rank, last device global rank
4.4. Communication Topology Scheduler
The communication topology scheduler employs a grid-search algorithm to discover the optimal parallelism configuration, tailored to the specifics of the current cluster, model, and input data.
When using Ring Attention, all GPUs form a single ring, which considerably limits the flexibility of adjusting the placement scheme. In contrast, WallFacer expands the tuning space with its unique multi-ring structure in two significant ways. First, the ring configuration in WallFacer is determined by the parallelism dimension $C\in[1,\sqrt{P}]$. Smaller $C$ results in less collective communication, a higher total communication volume, and more inter-node communication, while higher C leads to more communication volume during all-gather and reduce-scatter, which may exceed the overhead of the QKV matrix multiplication that we use for overlapping, but also highly reduce the overall communication volume.
Second, the additional dimension provided by WallFacer allows for two strategies in parallelism placement. One strategy is to keep the all-gather and reduce-scatter operations intra-node, which confines these processes within the same computing node. The other strategy is to keep the ring P2P communication intra-node, reducing latency and potentially enhancing performance. The choice between these strategies depends on the extent of overlapping achievable within the ring and the communication volume during collective communications.
Based on the above tuning space, we provide an auto-scheduler, grid-searching through all possible configurations for the one that provides the highest throughput, which can be described as:
(11) $\begin{split}Config=\arg\max_{C,placement}(Profile(C\in[1,\sqrt{P}],\\ placement\in[P2P\_intra,Collect\_intra])\end{split}$
The scheduler requires only a few iterations to profile the performance of automatically generated configurations. The time spent on this profiling is negligible compared to the entire training process, as it only needs to be performed once for each new computing cluster used for training.
Figure 6. A comparison of using naive and zigzag dataloader for 8 GPUs with attention parallel dimension of 2. The corresponding initialization can be found in Figure 5 with the same configuration. The improvement of efficiency from load-balancing increases with the number of GPUs.
4.5. Sequence Parallelism Dataloader
To accommodate both full mask and causal mask configurations for self-attention, we have implemented two distinct data loading schemes for load-balancing. For full masks, sequences are straightforwardly divided into equal sub-sequences and distributed among the devices. However, causal masks present a challenge due to the unbalanced computational load across GPUs; sub-sequences at the beginning of a sequence require significantly more computation than those at the end. To address this imbalance and achieve load equilibrium among GPUs, we modify the ZigZag scheme introduced by (Zhu et al., 2024), illustrated in Figure 6. The figure illustrates the simplest case of zigzag load-balancing. Notably, the effectiveness of this strategy improves as the number of GPUs increases. This improvement correlates with the expanding difference in computation volume between the first and the last token, which escalates as the sequence length extends. This approach ensures that the total workload on each GPU is balanced, eliminating the need for additional communication mechanisms like those employed in DistFlashAttention(Li et al., 2024).
4.6. WallFacer Implementation
WallFacer is written in PyTorch(Paszke et al., 2019) and uses the PyTorch torch.autograd.function and NCCL(NVIDIA, 2020) backend for forward and backward implementation. WallFacer also employs multiple techniques during runtime to improve its overall training efficiency.
Ingetrate Flash Attention. The WallFacer attention mechanism involves multiple iterations that loop over Keys and Values, with each iteration still using traditional self-attention with corresponding Query, Key, and Value (QKV). This approach enables WallFacer to incorporate flash attention effectively, extending its capability by preserving intermediate states across iterations. Additionally, WallFacer enhances the efficiency of the forward process with the help of torch JIT to fuse kernels aside from flash attention.
Overlap communication with computing. In WallFacer attention, P2P communication and self-attention computing are interleaved across iterations, each incurring considerable time. To mitigate this, WallFacer employs a double buffering technique to asynchronously execute communication and computing kernels, effectively overlapping these processes and enhancing GPU utilization.
Save recomputation with checkpoints. WallFacer adopts the checkpointing strategy introduced by DistFlashAttn(Li et al., 2024), placing checkpoints at the end of the self-attention phase rather than the FFN of each transformer layer. This checkpoint placement effectively obviates the need to recompute the self-attention forward process during the backward pass, avoiding redundant attention computation.
5. Evaluation
Figure 7. Throughput evaluation of Ring Attention and WallFacer on 32 GPUs from three different clusters. We place the performance of WallFacer with both C=2 and C=4 in the figure. The configurations are marked in the titles of the sub-figures. For instance, A100_8(3B, 63K) represents that the experiment is on machines with 8 Nvidia A100 GPUs in each node, the model used has three billion parameters, and the sequence length is 64k.
The computational resources we use in the experiments include a local computing cluster with eight nodes, each equipped with eight Nvidia H100 GPUs with 80GB of high-bandwidth memory, interconnected by NVLink and linked between nodes via InfiniBand with eight Mellanox ConnectX-7 controllers each. Additionally, we utilize Google Cloud Virtual Computing Machines in two setups: one with two machines containing 16 Nvidia A100 GPUs, and another with four machines containing eight GPUs each, both configurations using 40GB of high-bandwidth memory per GPU. All machines are connected within via NVLink and between by Ethernet.
We utilize two different model sizes for the two different GPU types. For H100 with 80GB HBM, we use a GPT model with approximately 7 billion parameters (referred to as 7B in the following text), which has 32 layers, 32 attention heads, and a hidden dimension of 4096. As for the A100 with 40 GB HBM, we use a smaller model with around 3 billion parameters (referred to as 3B), which has 12 layers, 16 attention heads, and a hidden dimension of 4096. During training, both models use bfloat16 precision and a batch size of 1 to accommodate longer input sequences.
In the evaluation section, we aim to answer three major questions:
  • How much improvement in throughput can WallFacer bring? Additionally, how adaptable is WallFacer to clusters with both good and poor inter-node connections?
  • Is the additional memory cost incurred by WallFacer acceptable considering the throughput improvement it offers?
  • How does WallFacer perform in scenarios of weak and strong scaling? Specifically, does it outperform Ring Attention when scaled to handle longer inputs?
Figure 8. The normalized relative memory cost of different configurations of WallFacer compared with Ring Attention on different clusters.
5.1. Throughput and Adaptability
Our first experiment aims to assess the performance of WallFacer and Ring Attention across different clusters with varying environments, testing the adaptability of both methods. There are several factors influencing the efficiency of ring-style attention computation:
Theoretical Computation-Communication Volume Ratio: Primarily determined by the sequence length used during training. Attention computation exhibits a computational complexity of $O(N^{2}\cdot H)$, whereas P2P communication complexity is $O(N\cdot H)$. Thus, the model configuration does not impact this ratio; only the sequence length does. A larger $N$ increases the computation-communication ratio, facilitating easier overlap of communication with computation. We evaluate varying sequence lengths to explore performance under different ratios.
Compute Capability and Connectivity of GPUs: The computing overhead, given a specific volume, affects the computation-communication overhead ratio. Higher compute capabilities make overlapping more challenging. We utilize two sets of GPUs in this evaluation: Nvidia A100 40GB and Nvidia H100 80GB, with the latter offering significantly higher theoretical tflops on bf16 computations. Connectivity is considered in two parts: intra-node and inter-node. Our clusters are equipped with NVLink, ensuring robust intra-node communication. For inter-node communication, while InfiniBand offers high bandwidth at a higher commercial cost, some clusters utilize Ethernet for connectivity. Specifically, our H100 nodes leverage InfiniBand with eight adapters per node for superior inter-node bandwidth, whereas the Google Cloud servers use Ethernet. The diversity in node configurations (8-GPU and 16-GPU nodes) allows us to assess adaptability across different topologies.
This evaluation not only highlights the inherent differences between the schemes but also tests their flexibility in various hardware settings.
The results of our evaluation are illustrated in Figure 7. We measure throughput in thousands of tokens per second. To better demonstrate how to select the optimal configuration of WallFacer under each condition, we included two configurations, Wall-2 and Wall-4, in the figure. We omit configurations with lower performance for clarity. As indicated in the figure, in all six settings, at least one configuration of WallFacer achieves higher throughput than Ring Attention, with performance improvements of 62.87%, 42.45%, 35.98%, 19.9%, 77.12%, and 34.62%. This advantage is primarily due to the additional parallel dimension that WallFacer introduces. Unlike Ring Attention, which requires inter-node P2P communication in each iteration, WallFacer’s P2P communication is mostly confined intra-node, except for initial data transfers. This experiment clearly demonstrates WallFacer’s superior performance across various environments.
Another observation is that the optimal configuration for WallFacer may vary depending on the environment, reflecting differences in the computation-communication ratio and the trade-offs between collective and P2P communication. For example, the best $C$ value for A100_16 is 2, while for A100_8, it is 4. This variation can be attributed to the topological differences: the 16-GPU-node cluster benefits less from reduced P2P communication volume, making a smaller $C$ preferable to decrease collective communication volume while maintaining P2P communication at an acceptable level. This finding underscores the importance of the Communication Topology Scheduler we provide, which helps users avoid the complex process of manually analyzing these factors.
In summary, WallFacer demonstrates superior efficiency and adaptability in all six cases, thanks to its flexible parallelism scheme.
5.2. Memory Consumption
Memory consumption in our experiments stems from both the model weights and activations, with additional costs for WallFacer arising from the duplication of QKV matrices prior to attention computation. To quantitatively assess the additional memory required by WallFacer, we monitored the maximum memory allocated by PyTorch (Paszke et al., 2019) during our experiments. It is important to note that memory fragmentation in PyTorch can impact memory allocation efficiency. To mitigate this, we limited the sequence lengths in our training to prevent PyTorch from triggering cuda_free when the allocated memory approaches the GPU’s limit, which would otherwise introduce significant overhead.
The results, displayed in Figure 8, reveal that for the configurations yielding the highest throughput, Ring Attention consumes between 7.9% and 30.79% less GPU memory than WallFacer. However, considering the substantial throughput gains provided by WallFacer, the additional memory usage is deemed acceptable. Additionally, evaluations for long sequences typically employ sequence lengths that are powers of two, and we observed that the maximum supported sequence lengths were not significantly impacted by this additional memory usage.
Moreover, in scenarios involving larger models with thirty billion (30B) or seventy billion (70B) parameters, the relative increase in memory consumption due to QKV duplication diminishes. This reduction is explained by equation 10, which shows that as the model size increases—owing to more parameters and additional Transformer layers—the proportion of memory consumed by model weights and activations grows, whereas the absolute extra memory incurred by QKV multiplication remains constant. This phenomenon is further evidenced by the fact that the extra memory ratio for experiments with the 7B model is significantly smaller than that for the 3B model.
Figure 9. Strong scaling experiments on Nvidia H100 GPUs with fixed sequence length of 128K.
5.3. Strong and Weak Scaling
In the scaling tests we carry out experiments for both strong and weak scaling. Strong Scaling maintains the scale of the problem that we are trying to solve while increasing the computing resource that we use to speed it up. So in this experiment, we fix the sequence length to 128K while increasing the number of GPUs from 8 to 64. As id depicted in Figure 9, WallFacer shows gradually more advantage over Ring Attention as we increase the number of GPUs. When we use 64 GPUs, WallFacer shows 60.71% higher throughput than Ring Attention. This can also explained by the computation-communication ratio. When scaled to more GPUs, the local sequence length on each GPU becomes smaller, and as explained the previous sections, makes it harder to overlap the P2P communication with attention computation. This is exactly the main advantage of WallFacer, reduction of total P2P communication volume. So WallFacer can give a promising performance when we are trying to speedup the training of a fixed amount of data with more GPUs.
Figure 10. Weak scaling on Nvidia H100 GPUs of sequence length from 64K to 512K
As for weak scaling, we can see from Figure 10 that WallFacer can still show gradually more advantage over Ring Attention, but not as much as that in strong scaling experiments. This is because the computation-communication ratio stays the same in weak scaling. But with more GPUs, Ring Attention will require a higher portion of inter-node communication, while WallFacer can remain most communication local.
In summary, WallFacer shows better scalability in both strong and weak scaling experiments, making it a better choice for large-scale Transformer model training.
6. Related Works
Attention Optimization. Traditional full attention mechanisms necessitate $O(n^{2})$ memory for storing the outputs of $QK^{T}$, leading to significant computational and memory demands. To address these challenges within the GPU, several approaches have been devised to reduce both memory and computational requirements. Memory-efficient attention(Rabe and Staats, 2022) introduces a straightforward algorithm that requires only $O(1)$ memory relative to the sequence length, with an extension for self-attention that needs only $O(\log n)$ memory. Flash Attention further minimizes I/O overhead and enhances overall efficiency. Additionally, optimization methods specifically tailored for inference, such as PagedAttention(Kwon et al., 2023), are also being developed to improve the efficiency of attention computations. In this work, we utilize Flash Attention within each iteration to reduce the computation overhead.
Long-Sequence Training Techniques. In recent years, numerous techniques have been developed for long-sequence training. Sequence Parallelism(Li et al., 2021) was initially introduced to enhance the efficiency of parallel long-sequence training. Ring Attention(Liu et al., 2023b) improved communication efficiency through memory-efficient methods(Rabe and Staats, 2022), supporting near-infinite sequence lengths. DeepSpeed Ulysses(Jacobs et al., 2023) employs attention head splitting to achieve high efficiency, though it is constrained by the number of heads. Megatron Sequence Parallelism focuses on reducing memory costs during Tensor Parallelism, while DistFlashAttention(Li et al., 2024) features a load-balance scheme and a novel gradient checkpoint method. Our work builds on these innovations, introducing a system that supports near-infinite sequence length and large-scale training with an efficient communication scheme.
Techniques for Distributed Model Training. Distributed model training encompasses two primary areas: 1) Memory Management: Various techniques aim to conserve GPU memory during distributed training, such as mixed precision training(Micikevicius et al., 2018) and the ZeRO series(Rajbhandari et al., 2020). Additionally, Zero-Offload(Ren et al., 2021) leverages CPU memory offloading. In this work, we implement ZeRO-2 to manage optimizer states and gradients efficiently. 2) Hybrid Parallelism: Frameworks like Megatron(Narayanan et al., 2021) and Colossal AI(Bian et al., 2021) integrate multiple forms of parallelism. There are various existing Parallelism techniques like Pipeline Parallelism(Huang et al., 2018; Fan et al., 2020; Li and Hoefler, 2021; Liu et al., 2023a) and Tensor Parallelism(Shoeybi et al., 2019), which can be combined with WallFacer Parallelism to facilitate large-scale training. We are also considering the integration of additional frameworks such as (Chen et al., 2024) to enhance overlapping capabilities in future implementations. In the future, we will examine hybrid strategies that combine WallFacer with other parallelisms.
7. Conclusion
WallFacer represents an advanced near-infinite-context Transformer model training system, featuring a communication-optimized multi-ring sequence parallelism scheme. In this study, we also conceptualize attention computation as a special case of the n-body problem, which introduces new perspectives and potential solutions for future challenges. Through comprehensive adaptability and scaling experiments, we demonstrate that our system not only achieves high efficiency across various training environments but also excels under both strong and weak scaling conditions. In an era increasingly demanding longer contexts for both natural language processing and computer vision, WallFacer is poised to make significant contributions to the industry and inspire innovative research in academia.

### v3

@@
-WallFacer: Guiding Transformer Model Training Out of the Long-Context Dark Forest with N-body Problem
-Abstract.
-In recent years, Transformer-based Large Language Models (LLMs) have garnered significant attention due to their exceptional performance across a variety of tasks.
-However, training these models on long sequences presents a substantial challenge in terms of efficiency and scalability.
-Current methods are constrained either by the number of attention heads, limiting scalability, or by excessive communication overheads.
-In this paper, we propose an insight that Attention Computation can be considered as a special case of n-body problem with direct interactions.
-Based on this concept, this paper introduces WallFacer, an efficient long-sequence training system with a novel multi-dimensional ring sequence parallelism, fostering an efficient communication paradigm and extra tuning space for communication arrangement.
-Through comprehensive experiments under diverse environments and model settings, we demonstrate that WallFacer significantly surpasses state-of-the-art method that supports near-infinite sequence length, achieving performance improvements of up to 77.12%.
-distributed deep learning, sequence parallelism, n-body simulation, large scale training, high performance computing
-copyright: noneconference: arxiv; Mar, 2025; ROTTERDAM
-1. Introduction
-Over the past decade, deep learning has made remarkable strides in diverse fields, including computer vision (CV) and natural language processing (NLP).
-Transformer-based (Vaswani et al., 2023) models, leveraging the groundbreaking attention mechanism introduced in 2018, have secured a dominant position due to their superior feature-capturing capabilities.
-As the technology has evolved, the ability to efficiently process long sequences has emerged as a pivotal challenge, capturing the attention of researchers and practitioners alike.
-Figure 1.
-Comparison of the theoretical peer-to-peer communication amount for Ring Attention and different configurations of WallFacer while training a llama-7B model(Touvron et al., 2023) on 64 GPUs with Adam optimizer(Kingma and Ba, 2017) and a batch size of 4.
-This focus on long-sequence training and inference is underscored by its critical importance in a variety of downstream tasks.
-For instance, in text summarization, the ability to handle extensive sequences is vital, as the content to be summarized can range from lengthy chapters to entire books (Koh et al., 2022; Beltagy et al., 2020).
-Similarly, chat-based applications, such as ChatGPT (Achiam et al., 2023), require the capacity to process extensive dialogue histories to ensure conversational consistency.
-There are also applications in other fields like video generation(Brooks et al., 2024; Peebles and Xie, 2023) and protein structure prediction(Jumper et al., 2021; Cheng et al., 2023).
-In summary, the demand for efficient long-sequence training and inference mechanisms spans a broad spectrum of applications, highlighting a critical area of need within the community.
+WallFacer: Harnessing Multi-dimensional Ring Parallelism for Efficient Long Sequence Model Training
+Abstract
+Training Transformer models on long sequences in a distributed setting poses significant challenges in terms of efficiency and scalability.
+Current methods are either constrained by the number of attention heads or excessive communication overheads.
+To address this problem, we propose WallFacer, a multi-dimensional distributed training system for long sequences, fostering an efficient communication paradigm and providing additional tuning flexibility for communication arrangements.
+Specifically, WallFacer introduces an extra parallel dimension to substantially reduce communication volume and avoid bandwidth bottlenecks.
+Through comprehensive experiments across diverse hardware environments and on both Natural Language Processing (NLP) and Computer Vision (CV) tasks, we demonstrate that our approach significantly surpasses state-of-the-art methods that support near-infinite sequence lengths, achieving performance improvements of up to 77.12% on GPT-style models and up to 114.33% on DiT (Diffusion Transformer) models.
+1 Introduction
+Over the past decade, Transformer[vaswani2023attention] models have made remarkable strides in diverse fields, including computer vision (CV) and natural language processing (NLP).
+As the technology has evolved, the ability to efficiently process long sequences with Transformer has emerged as a pivotal challenge.
+For instance, in text summarization, the ability to handle extensive sequences is vital, as the content to be summarized can range from lengthy chapters to entire books [Koh2022AnES, beltagy2020longformer].
+Similarly, chat-based applications, such as ChatGPT [openai2023gpt4], require the capacity to process extensive dialogue histories to ensure conversational consistency.
+There are also applications in other fields like video generation[Brooks_2024, peebles2023scalable] and protein structure prediction[Jumper2021HighlyAP, cheng2023fastfold].
+Figure 1: WallFacer-2 and WallFacer-4 theoretically save around 50% and 75% of total P2P communication volume for various sequence length.
 The long context in the above scenarios has introduced several challenges for model training and inference: 1) Efficiency and Adaptability.
-The challenge of efficiency is predominantly determined by computation and communication.
-How to maintain high efficiency in various environments and settings has drawn great attention from academia and industry.
+The challenge of efficiency predominantly lies in handling long sequences that require quadratic computations during attention, and in addressing the large amount of communication during distributed processing.
 2) Memory.
-As the size of the activation during attention is quadratically related to the length of the sequence, memory consumption increases rapidly with the growth of the sequence.
+Besides the major obstacle of storing the model weight and optimizer states, the activation has also exceeded the capacity of a single GPU and risen as a new memory challenge due to the extreme sequence length.
 3) Scalability.
-Large Language Models (LLMs) necessitate the use of thousands of GPUs for training, even with datasets of regular lengths.
-For long sequences, ensuring an acceptable scaling speedup rate is even more critical to reduce time and economic costs.
-Traditional parallelisms such as Data Parallelism(Hillis and Steele Jr, 1986; Shoeybi et al., 2019; Li et al., 2020; You et al., 2018), Tensor Parallelism(Shoeybi et al., 2019; Wang et al., 2022; Xu et al., 2021), and Pipeline Parallelism(Huang et al., 2018; Fan et al., 2020; Li and Hoefler, 2021; Liu et al., 2023a) have not been able to address the large memory requirement of extremely long sequences.
+Current Transformer models usually require thousands of GPUs for pre-training, even with datasets of regular lengths.
+For longer sequences, ensuring an acceptable scaling speedup rate with both the sequence length and the number of GPUs increasing is even more critical to reducing time and economic costs.
+Traditional parallelisms such as Data Parallelism[hillis1986data, shoeybi2019megatron, li2020taming, you2018imagenet], Tensor Parallelism[shoeybi2019megatron, wang2022tesseract, xu2021efficient], and Pipeline Parallelism[gpipe, dapple, li2021chimera, Liu2023Hanayo] distribute the model, input batch, and the optimizer states, but can not directly address the large memory requirement of extremely long sequences as the sequence length dimension remains unchanged.
 To break through this obstacle, Sequence Parallelism has been introduced, splitting the input on the sequence length dimension.
 Mainstream Sequence Parallelism schemes can generally be classified into two categories: those based on all-to-all communication, and those based on ring peer-to-peer communication.
-Methods like DeepSpeed Ulysses(Jacobs et al., 2023), which are based on all-to-all communication, offer efficiency but require the splitting of attention heads.
-Consequently, these methods are limited in scalability.
-On the other hand, peer-to-peer communication methods(Liu et al., 2023b; Li et al., 2024), such as Ring Attention(Liu et al., 2023b), do allow for infinite context lengths; however, they necessitate the transmission of complete keys and values across all GPUs, leading to significantly high communication loads.
-In summary, there remains a deficiency in communication-efficient methods that are capable of supporting infinite context lengths.
-In fact, we can draw valuable insights from other fields that have extensive experience in handling long sequences.
-In the field of n-body simulations, researchers have been dealing with vast numbers of particles, typically far exceeding the number of tokens addressed in current Transformer models.
-Through careful observation, we have identified that the core component of Transformers, the attention mechanism, closely resembles a special case of the n-body problem.
-Specifically, n-body problems with direct interactions focus on calculating the resultant force exerted by all other particles in the system on each individual particle.
-Similarly, in self-attention, the objective is to compute the attention score that each token in the sequence assigns to every other token.
-This similarity suggests that we can leverage methodologies from n-body problem research—refined over decades—to inform and enhance our approaches to tackling the relatively new challenges of long-context Transformer training.
-Inspired by the methodology of n-body communication optimization(Driscoll et al., 2013) and tailored to the Transformer architecture, we introduce WallFacer, a near-infinite-context Transformer training system with multi-ring sequence parallelism that incorporates an additional parallel dimension to the existing ring-style communication, fostering an efficient communication paradigm and extra tuning space for communication arrangement.
-With very little extra cost of memory, WallFacer parallelism significantly reduces the peer-to-peer communication amount, as is shown in figure 1.
-In summary, our paper presents these contributions:
-  • We conceptualize Attention computation as a novel instance of the traditional n-body problem, providing fresh insights into optimizing and parallelizing Attention computation.
-  • We introduce a near-infinite-context training system for Transformer models, featuring a groundbreaking multi-ring sequence parallelism scheme.
-This scheme adds an additional dimension of parallelism and significantly reduces peer-to-peer communication, all while maintaining a minimal memory footprint.
-  • We offer a straightforward method that allows users to select the most suitable parallelism scheme based on their specific needs, maximizing the utility of the available tuning space within our communication framework.
-  • We perform experiments on mainstream Transformer models, conducting performance and scaling tests across various computing clusters.
-Preliminary results indicate that our WallFacer system outperforms Ring Attention by up to 77.12%, showcasing its efficacy and scalability.
-2. Background
-2.1. Transformer and Long Sequence Training
-The key mechanism behind Transformer-based models is Attention(Vaswani et al., 2023), which captures the text feature by calculating the attention score between every two single tokens.
-A standard attention function is given as:
-(1) $Attention(Q,K,V)=softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V$
-where Q, K, and V are the Query, Value, and Key of input x with shape (B, N, H) in the case of self-attention, and $d_{k}$ represents the head dimension in commonly used multi-head attention.
-For convenience, we use the symbolic representations in Table 1.
-We can observe that in the original attention function, we need to store the intermediate value $QK^{T}$ of size $N^{2}$, which means that the memory capacity requirement and computation amount both grow in quadratic proportion with the sequence length.
-To solve the memory IO and capacity bottleneck, methods like flash-attention(Dao et al., 2022) have been proposed.
-Flash-attention is designed upon the idea of softmax decomposition (Kitaev et al., 2020; Milakov and Gimelshein, 2018; Rabe and Staats, 2022), enabling online softmax computation by storing some extra statistics.
-Flash-attention is thus capable of segmenting the Q, K, and V matrices into blocks, allowing the entire computation to be performed within a single kernel for each block sequentially.
-This strategy effectively circumvents redundant read and write operations to the High Bandwidth Memory (HBM).
-However, when the sequence length reaches millions, it becomes necessary to distribute the sequence across multiple GPUs.
-This distribution helps to reduce both the memory and computation demands on any single device.
-This strategy is also known as Sequence Parallelism.
-We will explore these methods in further detail in the following section.
-Table 1. Meanings of the symbols that are used in this paper
+Methods like DeepSpeed Ulysses[jacobs2023Ulysses], which are based on all-to-all communication, offer efficiency but require the splitting of attention heads.
+Consequently, these methods are limited in scalability and can not be scaled to more devices than the number of attention heads.
+On the other hand, peer-to-peer communication methods[liu2023ring, li2024distflashattn], such as Ring Attention[liu2023ring], do allow for near-infinite context lengths; however, they require the transmission of complete keys and values across all GPUs, leading to significantly high communication loads.
+In summary, there remains a deficiency in communication-efficient methods that are capable of supporting near-infinite context lengths.
+Table 1: Meanings of the symbols that are used in this paper
 $P$ The number of GPUs
@@
 $L$ The communication latency between GPUs
-2.2. Sequence Parallelism
-Sequence Parallelism was first conceptualized in a study published in 2021 (Li et al., 2021), which introduced the concept as a strategy to distribute computational sequences across multiple processing steps.
-In this paper, we adopt a broad definition of Sequence Parallelism to include any approach that partitions a sequence during specific computational phases.
+Inspired by the methodology of n-body communication optimization [Driscoll2013ACN], we introduce WallFacer, a novel near-infinite-context Transformer training system with multi-ring sequence parallelism that incorporates an additional parallel dimension into the existing ring-style communication.
+Specifically, instead of including all GPUs in a single parallel group as done in Ring Attention [liu2023ring], WallFacer groups the GPUs into teams and divides the peer-to-peer communication within these teams.
+This approach fosters an efficient communication paradigm and provides extra tuning flexibility for communication arrangements.
+With very little additional memory cost, WallFacer parallelism significantly reduces the peer-to-peer communication volume, as shown in Figure 1.
+Compared to previous works, WallFacer is not limited in supported sequence length by attention heads like DeepSpeed Ulysses[jacobs2023Ulysses] and Megatron Sequence Parallelism[korthikanti2022reducing], and also shows better communication efficiency and scalability than Ring Attention[liu2023ring].
+In summary, our paper presents these contributions:
+  • We introduce a near-infinite-context training system for Transformer models, featuring a groundbreaking multi-ring sequence parallelism scheme.
+This scheme adds an additional dimension of parallelism and significantly reduces peer-to-peer communication, all while maintaining a minimal memory footprint.
+  • We offer a straightforward grid-search method that allows users to select the most suitable parallelism scheme based on their specific needs, maximizing the utility of the available tuning space within our communication framework.
+  • We perform experiments on mainstream Transformer models, including GPT-style[gpt] and DiT-style[peebles2023scalablediffusionmodelstransformers], conducting performance and scaling tests across various computing clusters.
+Preliminary results indicate that our WallFacer system outperforms Ring Attention by up to 77.12% on the GPT model and up to 114.33% on the DiT model, showcasing its efficacy and scalability.
+2 Background and Motivation
+2.1 Transformer models and Long Sequence Training
+Transformer-based models have become the mainstream approach in both Natural Language Processing (NLP) and Computer Vision (CV).
+GPT-style models [gpt] lead the field in NLP tasks, generating tokens in an auto-regressive manner.
+Meanwhile, the Diffusion Transformer (DiT) [peebles2023scalablediffusionmodelstransformers], which employs pure Transformer blocks within traditional diffusion processes, has gained widespread use over the past year for image and video generation.
+The key mechanism behind these Transformer-based models is attention[vaswani2023attention], which captures the text feature by calculating the attention score between every two single tokens.
+A standard attention function is given as:
+$Attention(Q,K,V)=softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V$ (1)
+where $Q$, $K$, and $V$ are the query, value, and key of input x with shape ($B$, $N$, $H$) in the case of self-attention, and $d_{k}$ represents the head dimension in commonly used multi-head attention.
+For convenience, we use the symbolic representations in Table 1.
+We can observe that in the original attention function, we need to store the intermediate value $QK^{T}$ of size $N^{2}$, which means that the memory capacity requirement and computation amount both grow in quadratic proportion with the sequence length.
+To solve the memory IO and capacity bottleneck, methods like flash-attention[dao2022flashattention] have been proposed.
+Flash-attention is designed upon the idea of softmax decomposition [kitaev2020reformer, milakov2018online, rabe2022selfattention], enabling online softmax computation by storing some extra statistics.
+Flash-attention is thus capable of segmenting the $Q$, $K$, and $V$ matrices into blocks, allowing the entire computation to be performed within a single kernel for each block sequentially.
+This strategy effectively circumvents redundant read and write operations to the High Bandwidth Memory (HBM).
+However, the sequence length can reach hundreds of thousands, when dealing with multi-round chatting, or high-resolution long video generation.
+It then becomes necessary to distribute the sequence across multiple GPUs.
+This distribution helps to reduce both the memory and computation demands on any single device.
+This strategy is also known as sequence parallelism.
+We will explore these methods in further detail in the following section.
+2.2 Sequence Parallelism
+Sequence parallelism was first conceptualized in a study published in 2021 [Li2021SequencePL], which introduced the concept as a strategy to distribute computational sequences across multiple processing steps.
 Presently, Sequence parallelism can be divided into two main categories: attention-head-sharding-based and peer-to-peer-communication-based.
 The former involves distributing the attention heads of multi-head attention across multiple GPUs, whereas the latter resembles a distributed version of flash-attention, relying on peer-to-peer communication to transfer keys, values, and intermediate statistics.
-We will now explore the advantages and disadvantages of these methods in detail.
-2.2.1. Attention-head-Sharding-Based
+2.2.1 Attention-Head-Sharding-Based
 Here we introduce the two representative methods, DeepSpeed Ulysses and Megatron Sequence Parallelism.
 DeepSpeed Ulysses.
-DeepSpeed Ulysses(Jacobs et al., 2023) presents a strategy for training with long sequences by leveraging all-to-all collective communication, but is largely limited in its scalability.
-Functions outside the attention layer are partitioned along the sequence dimension using conventional methods.
-Within the attention block, Ulysses transitions from sequence parallelism to a method akin to tensor parallelism.
+DeepSpeed Ulysses[jacobs2023Ulysses] presents a strategy for training with long sequences by leveraging all-to-all collective communication, but is largely limited in its scalability.
+Within the attention block, Ulysses transitions from sequence parallelism to a method akin to tensor parallelism with two all-to-all communication.
 It divides the query, key, and value matrices across the attention heads, thereby preserving the original attention computation structure.
-This is facilitated by two sets of all-to-all communication that switch between sequence splitting and attention head spliting.
 The principal merit of Ulysses lies in its simplicity and ease of implementation.
 Nonetheless, its reliance on the number of attention heads for partitioning activations introduces the following limitations.
-1) Ulysses is limited in scalability, as it can only be expanded to as many GPUs as there are attention heads, which in turn restricts the maximum sequence length that can be processed.
-This limitation becomes particularly problematic when employing techniques like grouped-query attention (GQA) (Ainslie et al., 2023) or multi-query attention (MQA) (Shazeer, 2019), which further reduce the number of available attention heads for Keys and Values.
+1) Ulysses is limited in scalability, as it can only be expanded to as many GPUs as there are attention heads, which restricts the maximum sequence length that can be processed.
+This limitation becomes particularly problematic when employing techniques like grouped-query attention (GQA) [ainslie2023gqa] or multi-query attention (MQA) [shazeer2019fast], which further reduce the number of available attention heads for Keys and Values.
 2) Ulysses encounters load-balancing issues when the number of attention heads is not evenly divisible by the number of GPUs, complicating its deployment across diverse hardware configurations.
@@
 Megatron Sequence Parallelism.
-Megatron Sequence Parallelism(Korthikanti et al., 2022) shares its name with an earlier study(Li et al., 2021).
 Megatron Sequence Parallelism focuses on minimizing memory usage and reducing the necessity for activation recomputation.
@@
 This requires an additional all-gather operation and a further reduce-scatter operation during the forward propagation to alternate between Tensor Parallelism and Sequence Parallelism.
-Megatron Sequence Parallelism is not designed for training with exceedingly long sequences, as the Self-Attention and MLP layers continue to rely on Tensor Parallelism and must process the full sequence length.
+The Self-Attention and MLP layers continue to rely on Tensor Parallelism and must process the full sequence length.
 Like DeepSpeed Ulysses, it is also limited by the number of attention heads.
-2.2.2. Ring-peer-to-peer-communication-based
-The primary method in peer-to-peer-communication-based strategies is Ring Attention(Liu et al., 2023b).
+2.2.2 Ring-peer-to-peer-communication-based
+The primary method in peer-to-peer-communication-based strategies is Ring Attention[liu2023ring].
 Ring Attention.
-Introduced in 2023, Ring Attention(Liu et al., 2023b) innovatively partitions the sequence dimension and utilizes a ring-style peer-to-peer (P2P) communication pattern to transfer Keys and Values across all GPUs.
+Introduced in 2023, Ring Attention[liu2023ring] innovatively partitions the sequence dimension and utilizes a ring-style peer-to-peer (P2P) communication pattern to transfer Keys and Values across all GPUs.
 Each GPU receives the key and value matrices from the preceding rank, updates the local attention score, and then forwards them to the next rank, as is shown in Figure 2.
@@
 However, the requirement for $P$ steps of P2P communication renders this approach less efficient in environments with high-latency communication.
-A more detailed discussion on this topic is presented in the subsequent section.
-Figure 2. An example of Ring Attention Computation on 16 GPUs in two nodes. 16 rounds of ring P2P communication are needed to reach the final results.
-2.3. n-body Problem
-Before diving into the model formulation, it is essential to have a brief overview of the n-body simulation.
-N-body simulations(Hockney and Eastwood, 1989) play a crucial role in the study and modeling of dynamics within multi-body systems.
-Typically, the N in n-body represents the number of entities, often reaching up to $10^{12}$, far exceeding the processing capability of a single processor.
-Consequently, n-body simulations are predominantly conducted on supercomputing clusters.
-Specifically, n-body problems involving direct interaction refer to scenarios where only the direct forces between particles, such as gravitational forces, are considered.
-The design of n-body simulations encounters several significant challenges similar to those discussed in previous sections: 1) Efficiency: Continuous efforts are directed towards developing more efficient frameworks and algorithms to enhance both the computational and communicational aspects of n-body simulations.
-2) Memory: The number of particles in n-body problems typically surpasses those in natural language processing by orders of magnitude, presenting substantial challenges in managing memory on processors.
-3) Scalability: As n-body simulations are often executed on a large amount of computing nodes, scaling the computations to more processors is notably challenging.
-This backdrop of challenges in n-body simulations leads us to ponder whether the decades of accumulated knowledge in this field could be leveraged to inform and guide the training of Transformer models.
-3. Formulation: Transformer as N-body
-Through observation, we have discerned that attention computation can be viewed as a new instance of the n-body problem with direct interactions.
-Moreover, the forward and backward propagation processes in neural networks can be interpreted as distinct scenarios within this n-body framework.
-This insight suggests that the methodologies developed for n-body problems could also be effectively applied to the long-sequence training of Transformer models.
-3.1. Modeling of Attention as n-body
-The general n-body problem with direct interaction can be described as:
-(2) $B(x_{i})=\Sigma_{j=0}^{N}g(m_{0}(x_{j}),m_{1}(x_{i}))$
-Taking the cosmology gravity problem as an example, $B(x_{i})$ represents the total gravitational force experienced by an astronomical object x within the system, $m_{0}(x)$ and $m_{1}(x)$ both denote the mass of $x$ in this case, $g$ refers to the gravitational formula between two objects, and $\Sigma$ denotes the formula for the resultant force composition.
-Similarly, attention computation in neural networks can be expressed in an analogous format.
-Moreover, this formulation allows for distinct interpretations of the processes during forward and backward propagation: Forward Propagation.
-From the nature of matrix multiplication and formula 1, we can tell that the forward attention computation is independent for each $Query_{i}$, and each $Query_{i}$ needs to calculate score for the $Key$s and $Value$s of every token.
-In this way, we can rewrite formula 1 into formula 2 by following the Flash-Attention(Dao et al., 2022) style and having:
-(3) $\begin{split}m_{0}(x)&=W_{Q}x+B_{Q}=Q_{x}\\ m_{1}(x)&=W_{K}x+B_{K},\ W_{V}x+B_{V}=K_{x},V_{x}\\ g(m_{0}(x_{i}),m_{1}(x_{j}))&=softmax(\frac{Q_{xi}K^{T}_{xj}}{\sqrt{d_{k}}})V_{xj}\\ B(x_{i})&=\Sigma_{j=0}^{N}\ g(m_{0}(x_{i}),m_{1}(x_{j}))*e^{lse_{j}-lse_{new}}\end{split}$
-where the Ws and Bs refer to the weights and biases in the Linear functions, while lse is the LogSumExp (Nielsen and Sun, 2016) result calculated during $g$.
-Backward Propagation The primary distinction between forward and backward attention lies in the computational process during backward propagation.
-Specifically, to calculate the gradients for Keys and Values, it is necessary to iterate through all the Queries, and vice versa.
-This means that each token must interact with all three properties (Queries, Keys, Values) of every other token to compute its final gradients.
-We can analogize this to an n-body problem, where the interaction of forces is asymmetrical in two directions.
-Consequently, this requires conducting two rounds of the process as described in formula 2.
-In the first round, $x_{i}$ refers to $K_{i}$ and $V_{i}$ while in the second round, it refers to $Q_{i}$.
-In summary, we can easily interpret both the forward and backward attention calculation into a simple form of the n-body problem.
-Figure 3. An overview of the WallFacer Training System
-3.2. Methodology Transfer
-It is a recognized practice to harness the knowledge from traditional systems and high-performance computing to guide the design of machine learning systems.
-Now that we have modeled the attention calculation in Transformers as an n-body problem, we are poised to transfer a wealth of optimization methodologies from n-body simulation to Transformer training.
-This methodology transfer is deemed rational, as evidenced by existing techniques shared between these fields.
-For instance:
-Particle Decomposition: Commonly used to manage memory and computation constraints, this strategy involves distributing particles across multiple processors and integrating the results using the law of force composition.
-This method is widely implemented in models that employ online softmax, such as Flash-Attention(Dao et al., 2022).
-Hierarchical Methods: Hierarchical structures, such as the Multilevel Fast Multipole Method(Song and Chew, 1995; Lu and Chew, 1994) or tree-based processes like the Barnes–Hut simulation(Barnes and Hut, 1986), are used to approximate n-body calculations.
-These methods reduce the complexity from $O(n^{2})$ to $O(n\log n)$ or $O(n)$.
-Such approaches have been effectively applied in various fields, including computer vision, as seen in FASTERVIT(Hatamizadeh et al., 2024), etc., and in natural language processing tasks, as demonstrated in (Liu and Lapata, 2019) and so on.
-Cutoff: Interactions in many systems decay with distance, making it practical to set a cutoff distance and ignore interactions beyond this threshold.
-This principle is similarly utilized in approximate attention algorithms like Longformer(Beltagy et al., 2020), which employs a sliding window to calculate an approximate attention score for nearby tokens while excluding those outside the window.
-These examples validate the feasibility of guiding Transformer system design using n-body simulation techniques.
-Furthermore, we have identified additional aspects of n-body simulation that could inspire innovative approaches in training long-sequence Transformer models.
-In addition to the aforementioned techniques, we have explored a series of multi-dimensional algorithms used but not limited in n-body problems(Solomonik and Demmel, 2011; Aggarwal et al., 1990; Dekel et al., 1981; Agarwal et al., 1995; Gupta and Sadayappan, 1994) that optimize the communication scheme among processors while significantly reducing the total communication load, albeit at the expense of increased memory usage.
-The core idea involves leveraging an additional parallel dimension to partition the complete communication and computation tasks, distributing them across various devices.
-Inspired by these developments, we have conceived the concept of multi-ring sequence parallelism for our training system, which will be discussed in detail in the following section.
-4. WallFacer Training System
-4.1. Overview
-Built upon the inspiration from n-body simulation, the WallFacer training system is comprised of five main components.
+Figure 2: An example of Ring Attention Computation on 16 GPUs in two nodes. The Communication is largely limited by the inter-node bottleneck.
+Figure 3: An overview of the WallFacer Training System
+2.3 Motivation
+Through observation, we identify two main drawbacks of Ring Attention.
+First, the communication overhead is exceedingly high because every GPU in the system must send and receive keys and values for nearly the entire sequence length before completing the attention computation.
+Second, variations in bandwidth between and within computing nodes can cause communication bottlenecks.
+As illustrated in Figure 2, the bandwidths between GPUs 3 and 4 and between GPUs 11 and 12 are lower than those between other GPUs in the ring.
+Despite this, the system is forced to operate in a complete circle, which can result in unnecessary idle times for other GPUs.
+Inspired by techniques used in n-body simulations [Hockney_Eastwood_1989], we propose an approach that addresses both challenges simultaneously.
+In n-body simulations, such as those calculating direct forces between particles (e.g., gravitational forces), researchers discovered that communication overhead can be minimized by introducing additional dimensions of communication [Driscoll2013ACN].
+Similarly, we propose grouping GPUs into teams within the ring-style sequence parallelism framework and distributing communication tasks among team members.
+This approach reduces the overall communication volume and mitigates the bandwidth bottlenecks previously mentioned.
+Building upon this concept, we developed the WallFacer training system, which we will detail in the following section.
+3 WallFacer Training System
+3.1 Overview
+The WallFacer training system is comprised of five main components.
 At the core of the system is WallFacer Attention, which leverages multiple ring-style P2P (peer-to-peer) communication strategies to enhance the efficiency of distributed attention computation.
-The Dataloader is designed to organize tokens within each sub-sequence according to their mask (causal or full) and distribute them across different GPUs.
 The Communication Configuration Generator plays a critical role in initially assigning keys and values to their corresponding ranks.
 Meanwhile, the Communication Topology Scheduler outlines the placement of parallelism across various computing nodes.
+The Dataloader is designed to organize tokens within each sub-sequence according to their mask (causal or full) and distribute them across different GPUs.
 Finally, the WallFacer Runtime incorporates additional supporting techniques for the training process, such as gradient checkpointing.
 We will now explore the details of these components in depth.
-4.2. WallFacer Attention
-Figure 4. An example of a Transformer block with WallFacer Attention on one team of four devices out of 64 GPUs.
+3.2 WallFacer Attention
+Figure 4: An example of WallFacer Attention on 64 GPUs.
+Each team member forms a sub-ring with members with the same local rank from other teams in the same ring communication group, reducing the communication volume of each team member by 75%.
 As discussed in the previous section, a major limitation of Ring Attention is the extensive amount of peer-to-peer (P2P) communication required, which becomes problematic in environments with weak connections between computing nodes.
-Drawing inspiration from the multi-dimensional parallelism used in n-body simulations, we enhance the ring sequence parallelism by introducing an additional dimension.
+We enhance the ring sequence parallelism by introducing an additional dimension.
 This is achieved by duplicating the queries, keys, and values, thus dividing the communication tasks and distributing them within each team.
 This segmentation markedly improves computational efficiency and scalability.
-We will now delve into the specifics of the WallFacer Parallelism and provide a theoretical analysis demonstrating its advantages over Ring Attention.
-4.2.1. Training Process of WallFacer Parallelism
-In the WallFacer system, GPUs are grouped into Teams to coordinate computation and communication tasks more effectively, unlike the isolated operations in traditional settings.
+We will now go through the specifics of the WallFacer Parallelism and provide a theoretical analysis demonstrating its advantages over Ring Attention.
+3.2.1 Training Process of WallFacer Parallelism
+In the WallFacer system, GPUs are grouped into Teams to coordinate computation and communication tasks more efficiently, unlike the isolated operations in traditional settings.
 The queries, keys, and values are gathered within the team, so each member of the team holds the same activation of the whole team before attention.
@@
 When $C$ equals one, the algorithm falls back to Ring Attention.
-When C equals $\sqrt{P}$, the algorithm becomes a completely collective-communication-based one.
+When $C$ equals $\sqrt{P}$, the algorithm becomes a completely collective-communication-based one with no rings.
 When $1<C<\sqrt{P}$, it becomes a structure with multiple rings looping concurrently.
@@
 Algorithm 1 WallFacer Attention Block (Forward)
-0:  Input sequence x, Linear Function query, key, and value, attention parallelism size c, global rank r, global size gs, team process group pg
+0:  Input sequence x, Linear Function query, key, and value, attention parallelism size c, global rank r, global size gs, team process group pg, init send/recv target $\textbf{r}_{\mathrm{send}}$ and $\textbf{r}_{\mathrm{recv}}$
+1:  compute the gathered $\textbf{q}_{\mathrm{team}},\textbf{k}_{\mathrm{team}},\textbf{v}_{\mathrm{team}}=$AllGather_QKVmatmul(query, key, value, x, pg)
 […]
-2:  compute the initial rank to send $\textbf{r}_{send}=$$get\_init\_send(\textbf{r})$
-3:  compute the initial rank to receive from $\textbf{r}_{recv}=$$get\_init\_recv(\textbf{r})$
+3:  get the ring P2P target $\textbf{r}_{\mathrm{next}}$ and $\textbf{r}_{\mathrm{last}}$ withget_P2P_ranks(r, gs, c)
+4:  initialize attention score O, extra statistics lse to zero.
+5:  for $1\leq i\leq$world_size$/c^{2}$ do
+6:     wait for $\textbf{req}_{\mathrm{send}}$ and $\textbf{req}_{\mathrm{recv}}$
+7:     $\textbf{k}_{\mathrm{current}}=\textbf{k}_{\mathrm{next}}$, $\textbf{v}_{\mathrm{current}}=\textbf{v}_{\mathrm{next}}$
+8:     launch $\textbf{req}_{\mathrm{send}}$ to send $\textbf{k}_{\mathrm{current}}$ and $\textbf{v}_{\mathrm{current}}$ to $\textbf{r}_{\mathrm{next}}$, launch $\textbf{req}_{\mathrm{recv}}$ to receive $\textbf{k}_{\mathrm{next}}$ and $\textbf{v}_{\mathrm{next}}$ from $\textbf{r}_{\mathrm{last}}$
 […]
+10:  end for
 […]
-6:  initialize attention score O, extra statistics lse to zero.
-7:  for $1\leq i\leq world\_size/c^{2}$ do
-8:     wait for $\textbf{req}_{send}$ and $\textbf{req}_{recv}$
-9:     $\textbf{k}_{current}=\textbf{k}_{next}$, $\textbf{v}_{current}=\textbf{v}_{next}$
-10:     launch $\textbf{req}_{send}$ to send $\textbf{k}_{current}$ and $\textbf{v}_{current}$ to $\textbf{r}_{next}$, launch $\textbf{req}_{recv}$ to receive $\textbf{k}_{next}$ and $\textbf{v}_{next}$ from $\textbf{r}_{last}$
-[…]
-12:  end for
-[…]
-14:  return $\textbf{O}_{final}$
+12:  return $\textbf{O}_{\mathrm{final}}$
 Forward Propagation.
@@
 These are computed separately via matrix multiplication, followed immediately by the launch of the all-gather kernel, which gathers the above QKVs within the team, allowing for the overlap of up to two-thirds of the communication with computation.
-Once this phase is complete, each GPU within the team possesses the same Q, K, and Vs, each of a length of $\frac{CN}{P}$.
+Once this phase is complete, each GPU within the team possesses the same Q, K, and V, each of a length of $\frac{CN}{P}$.
 To distribute the communication and computation tasks among the team members, we divide the original workload based on four specific ranks assigned to each GPU by the Communication Configuration Generator.
@@
 Backward Propagation.
-The major distinction between backward and forward propagation, as outlined in section 3, is the inability to calculate Queries independently during the backward phase.
+The major distinction between backward and forward propagation is the inability to calculate queries independently during the backward phase.
 Unlike forward propagation, the backward phase requires the complete set of keys and values to calculate the gradient for queries, and vice versa.
@@
 The inner loop, however, handles the gradients for queries, which start initialized as zero and are circulated along the sub-rings together with the Queries themselves.
-During each iteration, the approach mirrors the backward computation method used in Flash Attention, where the updated gradient of the current Query shard is passed to the next GPU in the ring, while the gradients for Keys and Values are retained for subsequent Query shards.
+During each iteration, the approach mirrors the backward computation method used in flash-attention[dao2022flashattention], where the updated gradient of the current query shard is passed to the next GPU in the ring, while the gradients for keys and Values are retained for subsequent query shards.
 Initial communication ranks for sending and receiving are still configured by the Communication Configuration Generator.
 Additionally, there is an extra P2P communication required at the end of the process to send the query and its gradient back to their original location following the loop.
-4.2.2. Theoretical Analysis
-In this section, we will discuss two major questions: how much communication reduction can WallFacer bring compared with Ring Attention, and does this come with a high extra memory cost?
+3.2.2 Theoretical Analysis
+In this section, we will discuss two major questions: how much communication reduction WallFacer can bring compared with Ring Attention, and how much extra memory cost it comes with.
 During the analysis, we will employ a case study using the WallFacer system with an attention parallel size of $C$ = 4 on a llama-30B model, which consists of 64 layers.
@@
 For Ring Attention, the communication is primarily due to the ring P2P loop.
-Each iteration’s communication overhead can be calculated as follows:
-(4) $\frac{2BNH}{PW}+L$
-and as the total number of iterations done is $P$, the total communication overhead will be:
-(5) $\frac{2BNH}{W}+PL$
+As the total number of iterations done is $P$, the total communication overhead can be calculated as:
+$p(\frac{2BNH}{PW}+L)=\frac{2BNH}{W}+PL$ (2)
 and this overhead can be partially overlapped with the attention computation.
-For WallFacer, the communication overhead comes from collective and P2P both. The collective overhead for all-gather and reduce-scatter is:
-(6) $\frac{4BNH(C-1)}{PW}$
+For WallFacer, the communication overhead comes from both collective and P2P. The collective overhead for all-gather and reduce-scatter is:
+$\frac{4BNH(C-1)}{PW}$ (3)
 while the P2P communication can be similarly computed as:
-(7) $\frac{P}{C^{2}}(\frac{2CBNH}{PW}+L)=\frac{2BNH}{CW}+\frac{P}{C^{2}}L$
+$\frac{P}{C^{2}}(\frac{2CBNH}{PW}+L)=\frac{2BNH}{CW}+\frac{P}{C^{2}}L$ (4)
 The advantages of WallFacer over Ring Attention during the ring-P2P phase are evident in three main aspects: 1) Reduced Communication and Latency: Ring Attention requires C times more communication than WallFacer, significantly increasing the bandwidth requirement across the entire cluster.
@@
 In this section, we estimate the theoretical peak memory requirements necessary to store the model weights, activations, and optimizer states.
-Our implementation utilizes the Adam Optimizer (Kingma and Ba, 2017), bfloat16 precision, and Zero-2 optimization (Rajbhandari et al., 2020).
+Our implementation utilizes the Adam Optimizer [kingma2017adam], bfloat16 precision, and Zero-2 optimization [rajbhandari2020zero].
 Since the WallFacer architecture does not alter the model weights or the optimizer, we can assume that the memory costs associated with the model and the optimizer remain constant across both methods.
@@
 As for the activation, we refer to the size of one single activation of a sub-sequence on one GPU as
-(8) $A=\frac{B\times N\times H}{P}$
-As we use the checkpointing scheme from (Li et al., 2024), a model of $Y$ layers needs to save $Y+1$ activations as checkpoints.
+$A=\frac{B\times N\times H}{P}$ (5)
+As we use the checkpointing scheme from [li2024distflashattn], a model of $Y$ layers needs to save $Y+1$ activations as checkpoints.
 Now we calculate the approximate peak memory after Q, K, and V are already calculated and before the attention computation at the last layer of the whole model.
 For Ring Attention and WallFacer, the peak memories are:
-(9) $PM_{Ring}=M_{m+o}+(L+1)A+3A=M_{m+o}+(L+4)A$
-(10) $PM_{Wall}=M_{m+o}+(L+1)A+3CA=M_{m+o}+(L+3C+1)A$
+$PM_{Ring}=M_{m+o}+(Y+1)A+3A=M_{m+o}+(Y+4)A$ (6)
+$PM_{Wall}=M_{m+o}+(Y+1)A+3CA=M_{m+o}+(Y+3C+1)A$ (7)
 , where C is the WallFacer attention dimension.
@@
 In a word, the extra memory cost is acceptable as a tradeoff for the communication reduction.
-4.3. Communication Configuration Generator
-The Communication Topology Scheduler, as introduced in the previous section, plays a crucial role in the initial setup of Queries, Keys, and Values on GPUs and determining each GPU’s position within the sub-rings.
-Figure 5. An example of ring initialization process of 8GPUs and 4 sub-rings in the Communication Configuration Generator
+3.3 Communication Configuration Generator
+The Communication Configuration Generator, as introduced in the previous section, plays a crucial role in the initial setup of queries, keys, and values on GPUs and determining each GPU’s position within the sub-rings.
+Figure 5: An example of ring initialization process of 8GPUs and 4 sub-rings in the Communication Configuration Generator
 Initially, for the setup stage, it is essential to establish the sub-rings by rearranging the activation positions.
@@
 The number of teams within a team group can be calculated as $Team\ size=\frac{P}{C^{2}}$.
-We give the details of the determination of initial sending targets in the case of placing teams within same computing nodes in Algorithm 2, and the receiving source can also be calculated similarly.
+We give the details of the determination of initial sending targets in the case of placing teams within the same computing nodes in Algorithm 2, and the receiving source can also be calculated similarly.
 Algorithm 2 get_init_send()
@@
 7:  return next device global rank, last device global rank
-4.4. Communication Topology Scheduler
+3.4 Communication Topology Scheduler
 The communication topology scheduler employs a grid-search algorithm to discover the optimal parallelism configuration, tailored to the specifics of the current cluster, model, and input data.
@@
 First, the ring configuration in WallFacer is determined by the parallelism dimension $C\in[1,\sqrt{P}]$.
-Smaller $C$ results in less collective communication, a higher total communication volume, and more inter-node communication, while higher C leads to more communication volume during all-gather and reduce-scatter, which may exceed the overhead of the QKV matrix multiplication that we use for overlapping, but also highly reduce the overall communication volume.
+Smaller $C$ results in less collective communication, a higher total communication volume, and more inter-node communication, while higher C leads to more communication volume during all-gather and reduce-scatter, which may exceed the overhead of the QKV matrix multiplication that we use for overlapping, but can highly reduce the overall communication volume.
 Second, the additional dimension provided by WallFacer allows for two strategies in parallelism placement.
@@
 Based on the above tuning space, we provide an auto-scheduler, grid-searching through all possible configurations for the one that provides the highest throughput, which can be described as:
-(11) $\begin{split}Config=\arg\max_{C,placement}(Profile(C\in[1,\sqrt{P}],\\ placement\in[P2P\_intra,Collect\_intra])\end{split}$
+$\begin{split}Config=\arg\max_{C,placement}(Profile(C\in[1,\sqrt{P}],\\ placement\in[P2P\_intra,Collect\_intra])\end{split}$ (8)
 The scheduler requires only a few iterations to profile the performance of automatically generated configurations.
 The time spent on this profiling is negligible compared to the entire training process, as it only needs to be performed once for each new computing cluster used for training.
-Figure 6.
-A comparison of using naive and zigzag dataloader for 8 GPUs with attention parallel dimension of 2.
+Figure 6: A comparison between naive and zigzag dataloader for 8 GPUs with attention parallel dimension of 2.
 The corresponding initialization can be found in Figure 5 with the same configuration.
 The improvement of efficiency from load-balancing increases with the number of GPUs.
-4.5. Sequence Parallelism Dataloader
+3.5 Sequence Parallelism Dataloader
 To accommodate both full mask and causal mask configurations for self-attention, we have implemented two distinct data loading schemes for load-balancing.
-For full masks, sequences are straightforwardly divided into equal sub-sequences and distributed among the devices.
+For full masks, sequences are straightforwardly divided into sub-sequences of equal lengths and distributed among the devices.
 However, causal masks present a challenge due to the unbalanced computational load across GPUs; sub-sequences at the beginning of a sequence require significantly more computation than those at the end.
-To address this imbalance and achieve load equilibrium among GPUs, we modify the ZigZag scheme introduced by (Zhu et al., 2024), illustrated in Figure 6.
+To address this imbalance and achieve load equilibrium among GPUs, we modify the ZigZag scheme introduced by [zilin2024ringflashattention], illustrated in Figure 6.
 The figure illustrates the simplest case of zigzag load-balancing.
@@
 This improvement correlates with the expanding difference in computation volume between the first and the last token, which escalates as the sequence length extends.
-This approach ensures that the total workload on each GPU is balanced, eliminating the need for additional communication mechanisms like those employed in DistFlashAttention(Li et al., 2024).
-4.6. WallFacer Implementation
-WallFacer is written in PyTorch(Paszke et al., 2019) and uses the PyTorch torch.autograd.function and NCCL(NVIDIA, 2020) backend for forward and backward implementation.
+This approach ensures that the total workload on each GPU is balanced, eliminating the need for additional communication mechanisms like those employed in DistFlashAttention[li2024distflashattn].
+Figure 7: Throughput evaluation of Ring Attention and WallFacer on 32 GPUs from three different clusters.
+We place the performance of WallFacer with both C=2 and C=4 in the figure.
+The configurations are marked in the titles of the sub-figures.
+For instance, A100_40GB_8(1B, 512K) represents that the experiment is on machines with 8 Nvidia A100 40GB GPUs in each node, the model used has one billion parameters, and the sequence length is 512k.
+3.6 WallFacer Runtime
+WallFacer is written in PyTorch[pytorch] and uses the PyTorch torch.autograd.function and NCCL[nccl] backend for forward and backward implementation.
 WallFacer also employs multiple techniques during runtime to improve its overall training efficiency.
@@
 Save recomputation with checkpoints.
-WallFacer adopts the checkpointing strategy introduced by DistFlashAttn(Li et al., 2024), placing checkpoints at the end of the self-attention phase rather than the FFN of each transformer layer.
+WallFacer adopts the checkpointing strategy introduced by DistFlashAttn[li2024distflashattn], placing checkpoints at the end of the self-attention phase rather than the FFN of each transformer layer.
 This checkpoint placement effectively obviates the need to recompute the self-attention forward process during the backward pass, avoiding redundant attention computation.
-5. Evaluation
-Figure 7.
-Throughput evaluation of Ring Attention and WallFacer on 32 GPUs from three different clusters.
-We place the performance of WallFacer with both C=2 and C=4 in the figure.
-The configurations are marked in the titles of the sub-figures.
-For instance, A100_8(3B, 63K) represents that the experiment is on machines with 8 Nvidia A100 GPUs in each node, the model used has three billion parameters, and the sequence length is 64k.
-The computational resources we use in the experiments include a local computing cluster with eight nodes, each equipped with eight Nvidia H100 GPUs with 80GB of high-bandwidth memory, interconnected by NVLink and linked between nodes via InfiniBand with eight Mellanox ConnectX-7 controllers each.
-Additionally, we utilize Google Cloud Virtual Computing Machines in two setups: one with two machines containing 16 Nvidia A100 GPUs, and another with four machines containing eight GPUs each, both configurations using 40GB of high-bandwidth memory per GPU.
-All machines are connected within via NVLink and between by Ethernet.
-We utilize two different model sizes for the two different GPU types.
-For H100 with 80GB HBM, we use a GPT model with approximately 7 billion parameters (referred to as 7B in the following text), which has 32 layers, 32 attention heads, and a hidden dimension of 4096.
-As for the A100 with 40 GB HBM, we use a smaller model with around 3 billion parameters (referred to as 3B), which has 12 layers, 16 attention heads, and a hidden dimension of 4096.
+4 Evaluation
+The computational resources we use in the experiments include a local Nvidia H100 cluster with eight nodes and three Nvidia A100 clusters, as listed in table 2.
+We utilize two model types of total three settings, as listed in table 3.
+For the DiT(Diffusion Transformer) model, we use similar configurations as those in Stable Diffusion 3[esser2024sd3].
+We utilize the backbone Diffusion Transformer only, without other components like the text and image encoders.
 During training, both models use bfloat16 precision and a batch size of 1 to accommodate longer input sequences.
+Table 2: Cluster Configurations. All GPUs are connected with NVLink with computing nodes. #GPU represents local size times the number of nodes.
+GPU #GPUs Memory Inter-node
+Nvidia H100 8*8 80GB InfiniBand
+Nvidia A100 16*2 40GB 100Gbps Ethernet
+Nvidia A100 8*4 40GB 100Gbps Ethernet
+Nvidia A100 16*2 40GB 100Gbps Ethernet
+Table 3: Model Configurations
+Model Name #Heads #Layers Hidden Dim
+GPT 3B 12 16 4096
+GPT 7B 32 32 4096
+DiT 1B 24 24 1536
 In the evaluation section, we aim to answer three major questions:
@@
   • How does WallFacer perform in scenarios of weak and strong scaling? Specifically, does it outperform Ring Attention when scaled to handle longer inputs?
-Figure 8. The normalized relative memory cost of different configurations of WallFacer compared with Ring Attention on different clusters.
-5.1. Throughput and Adaptability
+Figure 8: The normalized relative memory cost of different configurations of WallFacer compared with Ring Attention on different clusters.
+4.1 Throughput and Adaptability
 Our first experiment aims to assess the performance of WallFacer and Ring Attention across different clusters with varying environments, testing the adaptability of both methods.
@@
 We omit configurations with lower performance for clarity.
-As indicated in the figure, in all six settings, at least one configuration of WallFacer achieves higher throughput than Ring Attention, with performance improvements of 62.87%, 42.45%, 35.98%, 19.9%, 77.12%, and 34.62%.
+As indicated in the figure, in all six settings, at least one configuration of WallFacer achieves higher throughput than Ring Attention, with performance improvements of 114.33%, 41.4%, 62.87%, 42.45%, 35.98%, 19.9%, 77.12%, and 34.62%.
 This advantage is primarily due to the additional parallel dimension that WallFacer introduces.
@@
 In summary, WallFacer demonstrates superior efficiency and adaptability in all six cases, thanks to its flexible parallelism scheme.
-5.2. Memory Consumption
+Table 4: Supported Sequence Length of Ring Attention and WallFacer on one Nvidia A100 80GB GPU.
+Supported Seq Len on one 80GB A100 GPU
+(K Tokens)
+Model Size Length Ring Attention WallFacer
+3B 128
+256
+512
+7B 128
+256
+512
+13B 128
+256
+512
+4.2 Memory Consumption
+Before evaluating the memory consumption of WallFacer, we first compared the supported sequence lengths of WallFacer with those reported in the RingAttention paper [liu2023ring].
+As shown in Table 4, despite the slightly higher memory requirements of WallFacer, it still supports sequence lengths commonly used in training tasks.
 Memory consumption in our experiments stems from both the model weights and activations, with additional costs for WallFacer arising from the duplication of QKV matrices prior to attention computation.
-To quantitatively assess the additional memory required by WallFacer, we monitored the maximum memory allocated by PyTorch (Paszke et al., 2019) during our experiments.
+To quantitatively assess the additional memory required by WallFacer, we monitored the maximum memory allocated by PyTorch [pytorch] during our experiments.
 It is important to note that memory fragmentation in PyTorch can impact memory allocation efficiency.
 To mitigate this, we limited the sequence lengths in our training to prevent PyTorch from triggering cuda_free when the allocated memory approaches the GPU’s limit, which would otherwise introduce significant overhead.
-The results, displayed in Figure 8, reveal that for the configurations yielding the highest throughput, Ring Attention consumes between 7.9% and 30.79% less GPU memory than WallFacer.
+The results, displayed in Figure 8, reveal that for the configurations yielding the highest throughput, Ring Attention consumes between 7.9% and 30.79% less GPU memory than WallFacer, except the DiT 512k Wall-4 setting.
 However, considering the substantial throughput gains provided by WallFacer, the additional memory usage is deemed acceptable.
-Additionally, evaluations for long sequences typically employ sequence lengths that are powers of two, and we observed that the maximum supported sequence lengths were not significantly impacted by this additional memory usage.
-Moreover, in scenarios involving larger models with thirty billion (30B) or seventy billion (70B) parameters, the relative increase in memory consumption due to QKV duplication diminishes.
+We observed that the maximum supported sequence lengths were not significantly impacted by this additional memory usage.
+Moreover, in scenarios involving larger models, the relative increase in memory consumption due to QKV duplication diminishes.
 This reduction is explained by equation 7, which shows that as the model size increases—owing to more parameters and additional Transformer layers—the proportion of memory consumed by model weights and activations grows, whereas the absolute extra memory incurred by QKV multiplication remains constant.
-This phenomenon is further evidenced by the fact that the extra memory ratio for experiments with the 7B model is significantly smaller than that for the 3B model.
-Figure 9. Strong scaling experiments on Nvidia H100 GPUs with fixed sequence length of 128K.
-5.3. Strong and Weak Scaling
+This phenomenon is further evidenced by the fact that the extra memory ratio for experiments with the 7B model is significantly smaller than that for the 3B and 1B model.
+4.3 Strong and Weak Scaling
 In the scaling tests we carry out experiments for both strong and weak scaling.
 Strong Scaling maintains the scale of the problem that we are trying to solve while increasing the computing resource that we use to speed it up.
-So in this experiment, we fix the sequence length to 128K while increasing the number of GPUs from 8 to 64.
-As id depicted in Figure 9, WallFacer shows gradually more advantage over Ring Attention as we increase the number of GPUs.
-When we use 64 GPUs, WallFacer shows 60.71% higher throughput than Ring Attention.
+So in this experiment, we fix the sequence length to 128K while increasing the number of GPUs from 8 to 64 for the GPT model and from 8 to 32 for the DiT model.
+As id depicted in Figure LABEL:Fig.strong, WallFacer shows gradually more advantage over Ring Attention as we increase the number of GPUs.
+When we use 64 GPUs, WallFacer shows 60.71% higher throughput than Ring Attention for the GPT model and 114.33% for the DiT model on 32 A100 GPUs.
 This can also explained by the computation-communication ratio.
@@
 So WallFacer can give a promising performance when we are trying to speedup the training of a fixed amount of data with more GPUs.
-Figure 10. Weak scaling on Nvidia H100 GPUs of sequence length from 64K to 512K
-As for weak scaling, we can see from Figure 10 that WallFacer can still show gradually more advantage over Ring Attention, but not as much as that in strong scaling experiments.
-This is because the computation-communication ratio stays the same in weak scaling.
-But with more GPUs, Ring Attention will require a higher portion of inter-node communication, while WallFacer can remain most communication local.
+In our weak scaling experiments using DiT models, we scale the sequence length from 128k to 512k while proportionally increasing the number of GPUs from 8 to 32.
+Assuming a scaling factor of $k$, both the sequence length and the number of GPUs increase by $k$, resulting in a total computation volume that increases by approximately $k^{2}$.
+Since the computational resources also increase by $k$, the theoretical overall throughput (tokens per second) should remain constant during this weak scaling setting.
+As shown in Figure LABEL:Fig.weak, both WallFacer and Ring Attention exhibit this expected behavior, with throughput remaining relatively constant.
+However, WallFacer consistently achieves higher throughput across different numbers of GPUs.
+This is because, although the local computation-to-communication ratio stays the same in weak scaling, Ring Attention requires a higher proportion of inter-node communication as the number of GPUs increases, whereas WallFacer keeps most communication local.
 In summary, WallFacer shows better scalability in both strong and weak scaling experiments, making it a better choice for large-scale Transformer model training.
-6. Related Works
+5 Discussion
+5.1 FlashAttention3 and Hopper GPUs
+In addition to the original FlashAttention [dao2022flashattention] used in our experiments, FlashAttention3 [shah2024flashattention3fastaccurateattention] has been introduced, specifically designed for Hopper and newer Nvidia GPUs.
+For FP16 precision, which is utilized in this paper, FlashAttention3 achieves a 1.5-2.0x speedup on Hopper GPUs.
+As discussed in Section 2, reducing attention computation overhead results in more P2P communication not being overlapped, further emphasizing the need to reduce communication volume.
+With the increasing adoption of Hopper GPUs, the significance of the WallFacer system will also grow.
+5.2 WallFacer and Other Parallelisms
+Model Parallelism As is well known, tensor parallelism shards activations by attention heads during attention computation, making it easily combinable with WallFacer with minimal effort.
+However, when combined with tensor parallelism, the need for attention heads can limit the scalability of head-based sequence parallel methods like Ulysses.
+Pipeline parallelism, on the other hand, divides the model across layers without altering the computation patterns within Transformer blocks, making WallFacer orthogonal to it.
+Other Sequence Parallelism WallFacer is orthogonal with other attention-head-sharding-based sequence parallelism approaches, such as DeepSpeed Ulysses [jacobs2023Ulysses].
+While Ulysses distributes attention heads across different devices, WallFacer can independently partition activations along the sequence length dimension.
+In future work, we can explore combining WallFacer with DeepSpeed Ulysses to expand the communication scheduling space, harnessing the scalability of WallFacer alongside the efficiency of Ulysses.
+In summary, WallFacer can be seamlessly integrated with other parallel training techniques, enabling the creation of a hybrid distributed training system.
+6 Related Work
 Attention Optimization.
@@
 To address these challenges within the GPU, several approaches have been devised to reduce both memory and computational requirements.
-Memory-efficient attention(Rabe and Staats, 2022) introduces a straightforward algorithm that requires only $O(1)$ memory relative to the sequence length, with an extension for self-attention that needs only $O(\log n)$ memory.
+Memory-efficient attention[rabe2022selfattention] introduces a straightforward algorithm that requires only $O(1)$ memory relative to the sequence length, with an extension for self-attention that needs only $O(\log n)$ memory.
 Flash Attention further minimizes I/O overhead and enhances overall efficiency.
-Additionally, optimization methods specifically tailored for inference, such as PagedAttention(Kwon et al., 2023), are also being developed to improve the efficiency of attention computations.
+Additionally, optimization methods specifically tailored for inference, such as PagedAttention[kwon2023efficient], are also being developed to improve the efficiency of attention computations.
 In this work, we utilize Flash Attention within each iteration to reduce the computation overhead.
 Long-Sequence Training Techniques.
-In recent years, numerous techniques have been developed for long-sequence training.
-Sequence Parallelism(Li et al., 2021) was initially introduced to enhance the efficiency of parallel long-sequence training.
-Ring Attention(Liu et al., 2023b) improved communication efficiency through memory-efficient methods(Rabe and Staats, 2022), supporting near-infinite sequence lengths.
-DeepSpeed Ulysses(Jacobs et al., 2023) employs attention head splitting to achieve high efficiency, though it is constrained by the number of heads.
-Megatron Sequence Parallelism focuses on reducing memory costs during Tensor Parallelism, while DistFlashAttention(Li et al., 2024) features a load-balance scheme and a novel gradient checkpoint method.
-Our work builds on these innovations, introducing a system that supports near-infinite sequence length and large-scale training with an efficient communication scheme.
+Sequence Parallelism[Li2021SequencePL] was initially introduced to enhance the efficiency of parallel long-sequence training.
+Ring Attention[liu2023ring] improved communication efficiency through memory-efficient methods[rabe2022selfattention], supporting near-infinite sequence lengths.
+DeepSpeed Ulysses[jacobs2023Ulysses] employs attention head splitting to achieve high efficiency, though it is constrained by the number of heads.
+Megatron Sequence Parallelism focuses on reducing memory costs during Tensor Parallelism, while DistFlashAttention[li2024distflashattn] features a load-balance scheme and a novel gradient checkpoint method.
+Our work builds on these innovations, introducing a system that supports large-scale training with an efficient communication scheme.
 Techniques for Distributed Model Training.
-Distributed model training encompasses two primary areas: 1) Memory Management: Various techniques aim to conserve GPU memory during distributed training, such as mixed precision training(Micikevicius et al., 2018) and the ZeRO series(Rajbhandari et al., 2020).
-Additionally, Zero-Offload(Ren et al., 2021) leverages CPU memory offloading.
+Distributed model training encompasses two primary areas: 1) Memory Management: Various techniques aim to conserve GPU memory during distributed training, such as mixed precision training[micikevicius2018mixed] and the ZeRO series[rajbhandari2020zero].
 In this work, we implement ZeRO-2 to manage optimizer states and gradients efficiently.
-2) Hybrid Parallelism: Frameworks like Megatron(Narayanan et al., 2021) and Colossal AI(Bian et al., 2021) integrate multiple forms of parallelism.
-There are various existing Parallelism techniques like Pipeline Parallelism(Huang et al., 2018; Fan et al., 2020; Li and Hoefler, 2021; Liu et al., 2023a) and Tensor Parallelism(Shoeybi et al., 2019), which can be combined with WallFacer Parallelism to facilitate large-scale training.
-We are also considering the integration of additional frameworks such as (Chen et al., 2024) to enhance overlapping capabilities in future implementations.
-In the future, we will examine hybrid strategies that combine WallFacer with other parallelisms.
-7. Conclusion
+2) Hybrid Parallelism: Frameworks like Megatron[megatron] and Colossal AI[bian2021colossal] integrate multiple forms of parallelism.
+There are various existing Parallelism techniques like Pipeline Parallelism[gpipe, dapple, li2021chimera, Liu2023Hanayo] and Tensor Parallelism[shoeybi2019megatron], which can be combined with WallFacer Parallelism to facilitate large-scale training.
+We are also considering the integration of additional frameworks such as [chang2024centauri] to enhance overlapping capabilities in future implementations.
+7 Conclusion
 WallFacer represents an advanced near-infinite-context Transformer model training system, featuring a communication-optimized multi-ring sequence parallelism scheme.
-In this study, we also conceptualize attention computation as a special case of the n-body problem, which introduces new perspectives and potential solutions for future challenges.
-Through comprehensive adaptability and scaling experiments, we demonstrate that our system not only achieves high efficiency across various training environments but also excels under both strong and weak scaling conditions.
-In an era increasingly demanding longer contexts for both natural language processing and computer vision, WallFacer is poised to make significant contributions to the industry and inspire innovative research in academia.
+Through comprehensive adaptability and scaling experiments, we demonstrate that our system not only achieves high efficiency across various training environments but also excels under both strong and weak scaling conditions for both Computer Vision(CV) and Natral Language Processing(NLP) models.
+In an era increasingly demanding longer contexts for both NLP and CV, WallFacer is poised to make significant contributions to the industry and inspire innovative research in academia.
+[…]

### v4

@@
-WallFacer: Harnessing Multi-dimensional Ring Parallelism for Efficient Long Sequence Model Training
+StarTrail: Concentric Ring Sequence Parallelism for Efficient Near-Infinite-Context Transformer Model Training
 Abstract
@@
 Current methods are either constrained by the number of attention heads or excessive communication overheads.
-To address this problem, we propose WallFacer, a multi-dimensional distributed training system for long sequences, fostering an efficient communication paradigm and providing additional tuning flexibility for communication arrangements.
-Specifically, WallFacer introduces an extra parallel dimension to substantially reduce communication volume and avoid bandwidth bottlenecks.
-Through comprehensive experiments across diverse hardware environments and on both Natural Language Processing (NLP) and Computer Vision (CV) tasks, we demonstrate that our approach significantly surpasses state-of-the-art methods that support near-infinite sequence lengths, achieving performance improvements of up to 77.12% on GPT-style models and up to 114.33% on DiT (Diffusion Transformer) models.
+To address this problem, we propose StarTrail, a multi-dimensional concentric distributed training system for long sequences, fostering an efficient communication paradigm and providing additional tuning flexibility for communication arrangements.
+Specifically, StarTrail introduces an extra parallel dimension and divides the peer-to-peer communication into sub-rings to substantially reduce communication volume and avoid bandwidth bottlenecks.
+Through comprehensive experiments across diverse hardware environments and on both Natural Language Processing (NLP) and Computer Vision (CV) tasks, we demonstrate that our approach significantly surpasses state-of-the-art methods that support Long sequence lengths, achieving performance improvements of up to 77.12% on GPT-style models and up to 114.33% on DiT (Diffusion Transformer) models without affecting the computation results.
 1 Introduction
-Over the past decade, Transformer[vaswani2023attention] models have made remarkable strides in diverse fields, including computer vision (CV) and natural language processing (NLP).
+Over the past decade, Transformervaswani2023attention models have made remarkable strides in diverse fields, including computer vision (CV) and natural language processing (NLP).
 As the technology has evolved, the ability to efficiently process long sequences with Transformer has emerged as a pivotal challenge.
-For instance, in text summarization, the ability to handle extensive sequences is vital, as the content to be summarized can range from lengthy chapters to entire books [Koh2022AnES, beltagy2020longformer].
-Similarly, chat-based applications, such as ChatGPT [openai2023gpt4], require the capacity to process extensive dialogue histories to ensure conversational consistency.
-There are also applications in other fields like video generation[Brooks_2024, peebles2023scalable] and protein structure prediction[Jumper2021HighlyAP, cheng2023fastfold].
-Figure 1: WallFacer-2 and WallFacer-4 theoretically save around 50% and 75% of total P2P communication volume for various sequence length.
+For instance, in text summarization, the ability to handle extensive sequences is vital, as the content to be summarized can range from lengthy chapters to entire books Koh2022AnES ; beltagy2020longformer .
+Similarly, chat-based applications, such as ChatGPT openai2023gpt4 , require the capacity to process extensive dialogue histories to ensure conversational consistency.
+There are also applications in other fields like video generationBrooks_2024 ; peebles2023scalable and protein structure predictionJumper2021HighlyAP ; cheng2023fastfold .
 The long context in the above scenarios has introduced several challenges for model training and inference: 1) Efficiency and Adaptability.
@@
 For longer sequences, ensuring an acceptable scaling speedup rate with both the sequence length and the number of GPUs increasing is even more critical to reducing time and economic costs.
-Traditional parallelisms such as Data Parallelism[hillis1986data, shoeybi2019megatron, li2020taming, you2018imagenet], Tensor Parallelism[shoeybi2019megatron, wang2022tesseract, xu2021efficient], and Pipeline Parallelism[gpipe, dapple, li2021chimera, Liu2023Hanayo] distribute the model, input batch, and the optimizer states, but can not directly address the large memory requirement of extremely long sequences as the sequence length dimension remains unchanged.
+Figure 1: StarTrail-2 and StarTrail-4 theoretically save around 50% and 75% of total P2P communication volume for various sequence length and achieves up to 2x speedup in end-to-end training
+Traditional parallelisms such as Data Parallelismhillis1986data ; shoeybi2019megatron ; li2020taming ; you2018imagenet , Tensor Parallelismshoeybi2019megatron ; wang2022tesseract ; xu2021efficient , and Pipeline Parallelismgpipe ; dapple ; li2021chimera ; Liu2023Hanayo distribute the model, input batch, and the optimizer states, but can not directly address the large memory requirement of extremely long sequences as the sequence length dimension remains unchanged.
 To break through this obstacle, Sequence Parallelism has been introduced, splitting the input on the sequence length dimension.
-Mainstream Sequence Parallelism schemes can generally be classified into two categories: those based on all-to-all communication, and those based on ring peer-to-peer communication.
-Methods like DeepSpeed Ulysses[jacobs2023Ulysses], which are based on all-to-all communication, offer efficiency but require the splitting of attention heads.
+Mainstream Sequence Parallelism schemes can generally be classified into two categories, and are usually combined fang2024uspunifiedsequenceparallelism to complement each other’s drawbacks.
+Methods like DeepSpeed Ulyssesjacobs2023Ulysses , which are based on all-to-all communication, offer efficiency but require the splitting of attention heads.
 Consequently, these methods are limited in scalability and can not be scaled to more devices than the number of attention heads.
-On the other hand, peer-to-peer communication methods[liu2023ring, li2024distflashattn], such as Ring Attention[liu2023ring], do allow for near-infinite context lengths; however, they require the transmission of complete keys and values across all GPUs, leading to significantly high communication loads.
+On the other hand, peer-to-peer communication methodsliu2023ring ; li2024distflashattn , such as Ring Attentionliu2023ring , do allow for near-infinite context lengths; however, they require the transmission of complete keys and values across all GPUs, leading to significantly high communication loads.
 In summary, there remains a deficiency in communication-efficient methods that are capable of supporting near-infinite context lengths.
-Table 1: Meanings of the symbols that are used in this paper
-$P$ The number of GPUs
-$C$ The parallel size of WallFacer (attention parallel size)
-$H$ The hidden dimension size of the Transformer blocks
-$N$ The total number of tokens within the whole sequence
-$B$ The training batch size
-$W$ The communication bandwidth between GPUs
-$L$ The communication latency between GPUs
-Inspired by the methodology of n-body communication optimization [Driscoll2013ACN], we introduce WallFacer, a novel near-infinite-context Transformer training system with multi-ring sequence parallelism that incorporates an additional parallel dimension into the existing ring-style communication.
-Specifically, instead of including all GPUs in a single parallel group as done in Ring Attention [liu2023ring], WallFacer groups the GPUs into teams and divides the peer-to-peer communication within these teams.
+In this paper, we focus on solving the communication inefficiency of ring-style sequence parallelism.
+To solve these challenges, we introduce StarTrail, a novel near-infinite-context Transformer training system with concentric multi-ring sequence parallelism that incorporates an additional parallel dimension into the existing ring-style communication.
+Specifically, instead of including all GPUs in a single parallel group as done in Ring Attention liu2023ring , StarTrail groups the GPUs into teams and divides the peer-to-peer communication within these teams.
 This approach fosters an efficient communication paradigm and provides extra tuning flexibility for communication arrangements.
-With very little additional memory cost, WallFacer parallelism significantly reduces the peer-to-peer communication volume, as shown in Figure 1.
-Compared to previous works, WallFacer is not limited in supported sequence length by attention heads like DeepSpeed Ulysses[jacobs2023Ulysses] and Megatron Sequence Parallelism[korthikanti2022reducing], and also shows better communication efficiency and scalability than Ring Attention[liu2023ring].
-In summary, our paper presents these contributions:
-  • We introduce a near-infinite-context training system for Transformer models, featuring a groundbreaking multi-ring sequence parallelism scheme.
-This scheme adds an additional dimension of parallelism and significantly reduces peer-to-peer communication, all while maintaining a minimal memory footprint.
-  • We offer a straightforward grid-search method that allows users to select the most suitable parallelism scheme based on their specific needs, maximizing the utility of the available tuning space within our communication framework.
-  • We perform experiments on mainstream Transformer models, including GPT-style[gpt] and DiT-style[peebles2023scalablediffusionmodelstransformers], conducting performance and scaling tests across various computing clusters.
-Preliminary results indicate that our WallFacer system outperforms Ring Attention by up to 77.12% on the GPT model and up to 114.33% on the DiT model, showcasing its efficacy and scalability.
-2 Background and Motivation
-2.1 Transformer models and Long Sequence Training
-Transformer-based models have become the mainstream approach in both Natural Language Processing (NLP) and Computer Vision (CV).
-GPT-style models [gpt] lead the field in NLP tasks, generating tokens in an auto-regressive manner.
-Meanwhile, the Diffusion Transformer (DiT) [peebles2023scalablediffusionmodelstransformers], which employs pure Transformer blocks within traditional diffusion processes, has gained widespread use over the past year for image and video generation.
-The key mechanism behind these Transformer-based models is attention[vaswani2023attention], which captures the text feature by calculating the attention score between every two single tokens.
-A standard attention function is given as:
-$Attention(Q,K,V)=softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V$ (1)
-where $Q$, $K$, and $V$ are the query, value, and key of input x with shape ($B$, $N$, $H$) in the case of self-attention, and $d_{k}$ represents the head dimension in commonly used multi-head attention.
-For convenience, we use the symbolic representations in Table 1.
-We can observe that in the original attention function, we need to store the intermediate value $QK^{T}$ of size $N^{2}$, which means that the memory capacity requirement and computation amount both grow in quadratic proportion with the sequence length.
-To solve the memory IO and capacity bottleneck, methods like flash-attention[dao2022flashattention] have been proposed.
-Flash-attention is designed upon the idea of softmax decomposition [kitaev2020reformer, milakov2018online, rabe2022selfattention], enabling online softmax computation by storing some extra statistics.
-Flash-attention is thus capable of segmenting the $Q$, $K$, and $V$ matrices into blocks, allowing the entire computation to be performed within a single kernel for each block sequentially.
-This strategy effectively circumvents redundant read and write operations to the High Bandwidth Memory (HBM).
+With very little additional memory cost, StarTrail parallelism significantly reduces the peer-to-peer communication volume, as shown in Figure 1.
+Compared to previous works, StarTrail is not limited in supported sequence length by attention heads like DeepSpeed Ulyssesjacobs2023Ulysses and Megatron Sequence Parallelismkorthikanti2022reducing , and also shows better communication efficiency and scalability than Ring Attentionliu2023ring .
+We perform experiments on mainstream Transformer models, including GPT-stylegpt and DiT-stylepeebles2023scalablediffusionmodelstransformers , conducting performance and scaling tests across various computing clusters.
+Experiment results indicate that our StarTrail system outperforms Ring Attention by up to 77.12% on the GPT model and up to 114.33% on the DiT model, showcasing its efficiency and scalability.
+2 Background and Related Works
+2.1 Long Sequence Training and Sequence Parallelism
+The key mechanism behind these Transformer-based models is attentionvaswani2023attention , which captures the text feature by calculating the attention score between every two single tokens.
 However, the sequence length can reach hundreds of thousands, when dealing with multi-round chatting, or high-resolution long video generation.
@@
 This strategy is also known as sequence parallelism.
-We will explore these methods in further detail in the following section.
-2.2 Sequence Parallelism
-Sequence parallelism was first conceptualized in a study published in 2021 [Li2021SequencePL], which introduced the concept as a strategy to distribute computational sequences across multiple processing steps.
 Presently, Sequence parallelism can be divided into two main categories: attention-head-sharding-based and peer-to-peer-communication-based.
-The former involves distributing the attention heads of multi-head attention across multiple GPUs, whereas the latter resembles a distributed version of flash-attention, relying on peer-to-peer communication to transfer keys, values, and intermediate statistics.
-2.2.1 Attention-Head-Sharding-Based
-Here we introduce the two representative methods, DeepSpeed Ulysses and Megatron Sequence Parallelism.
-DeepSpeed Ulysses.
-DeepSpeed Ulysses[jacobs2023Ulysses] presents a strategy for training with long sequences by leveraging all-to-all collective communication, but is largely limited in its scalability.
-Within the attention block, Ulysses transitions from sequence parallelism to a method akin to tensor parallelism with two all-to-all communication.
-It divides the query, key, and value matrices across the attention heads, thereby preserving the original attention computation structure.
-The principal merit of Ulysses lies in its simplicity and ease of implementation.
-Nonetheless, its reliance on the number of attention heads for partitioning activations introduces the following limitations.
-1) Ulysses is limited in scalability, as it can only be expanded to as many GPUs as there are attention heads, which restricts the maximum sequence length that can be processed.
-This limitation becomes particularly problematic when employing techniques like grouped-query attention (GQA) [ainslie2023gqa] or multi-query attention (MQA) [shazeer2019fast], which further reduce the number of available attention heads for Keys and Values.
-2) Ulysses encounters load-balancing issues when the number of attention heads is not evenly divisible by the number of GPUs, complicating its deployment across diverse hardware configurations.
-3) The need for head-splitting makes Ulysses hard to combine with Tensor Parallelism.
-Megatron Sequence Parallelism.
-Megatron Sequence Parallelism focuses on minimizing memory usage and reducing the necessity for activation recomputation.
-The training approach is based on Tensor Parallelism (TP).
-While TP is employed in the Linear and Self-Attention blocks, operations such as the LayerNorm and Dropout functions are not distributed, leading to replications within the tensor parallel group.
-This requires an additional all-gather operation and a further reduce-scatter operation during the forward propagation to alternate between Tensor Parallelism and Sequence Parallelism.
-The Self-Attention and MLP layers continue to rely on Tensor Parallelism and must process the full sequence length.
-Like DeepSpeed Ulysses, it is also limited by the number of attention heads.
+The former involves distributing the attention heads of multi-head attention across multiple GPUs, whereas the latter resembles a distributed version of FlashAttention, relying on peer-to-peer communication to transfer keys, values, and intermediate statistics.
+Figure 2: An example of Ring Attention Computation on 16 GPUs in two nodes. The Communication is largely limited by the inter-node bottleneck.
 2.1.1 Ring-peer-to-peer-communication-based
-The primary method in peer-to-peer-communication-based strategies is Ring Attention[liu2023ring].
-Ring Attention.
-Introduced in 2023, Ring Attention[liu2023ring] innovatively partitions the sequence dimension and utilizes a ring-style peer-to-peer (P2P) communication pattern to transfer Keys and Values across all GPUs.
+The primary method in peer-to-peer-communication-based strategies is Ring Attentionliu2023ring , which is also the main baseline of this work.
+Introduced in 2023, Ring Attentionliu2023ring innovatively partitions the sequence dimension and utilizes a ring-style peer-to-peer (P2P) communication pattern to transfer Keys and Values across all GPUs.
 Each GPU receives the key and value matrices from the preceding rank, updates the local attention score, and then forwards them to the next rank, as is shown in Figure 2.
@@
 Thus, it potentially supports infinite context, provided sufficient computing resources are available.
-However, the requirement for $P$ steps of P2P communication renders this approach less efficient in environments with high-latency communication.
-Figure 2: An example of Ring Attention Computation on 16 GPUs in two nodes. The Communication is largely limited by the inter-node bottleneck.
-Figure 3: An overview of the WallFacer Training System
+However, the requirement for the same number of rounds of P2P communication as the number of GPUs renders this approach less efficient in environments with high-latency communication.
+Figure 3: An overview of the StarTrail Training System
+2.1.2 Attention-Head-Sharding-Based
+There are two representative methods, DeepSpeed-Ulysses and Megatron Sequence Parallelism in this category.
+DeepSpeed-Ulyssesjacobs2023Ulysses transitions from sequence parallelism to a method akin to tensor parallelism with two all-to-all communication.
+It divides the query, key, and value matrices across the attention heads, thereby preserving the original attention computation structure.
+Megatron Sequence Parallelismkorthikanti2022reducing focuses on minimizing memory usage and reducing the necessity for activation recomputation rather than efficiency.
+The two methods both rely on the number of attention heads to split the sequence, thus limited in scalability, especially when employing techniques like grouped-query attention (GQA) ainslie2023gqa or multi-query attention (MQA) shazeer2019fast .
+As these two sequence parallelism methods are orthogonal to the ring-based method, they are usually combined with ring attention to enable longer sequences.
+In this paper, we focus on optimizing ring-style sequence parallelism, noting that integrating it with DeepSpeed-Ulysses Parallelism does not interfere with the underlying ring process.
+3 StarTrail Training System
+Figure 4: An example of StarTrail Attention on 64 GPUs.
+Each team member forms a sub-ring with members with the same local rank from other teams in the same ring communication group, reducing the communication volume of each team member by 75%.
 3.1 Motivation
@@
 Despite this, the system is forced to operate in a complete circle, which can result in unnecessary idle times for other GPUs.
-Inspired by techniques used in n-body simulations [Hockney_Eastwood_1989], we propose an approach that addresses both challenges simultaneously.
-In n-body simulations, such as those calculating direct forces between particles (e.g., gravitational forces), researchers discovered that communication overhead can be minimized by introducing additional dimensions of communication [Driscoll2013ACN].
-Similarly, we propose grouping GPUs into teams within the ring-style sequence parallelism framework and distributing communication tasks among team members.
-This approach reduces the overall communication volume and mitigates the bandwidth bottlenecks previously mentioned.
-Building upon this concept, we developed the WallFacer training system, which we will detail in the following section.
-3 WallFacer Training System
-3.1 Overview
-The WallFacer training system is comprised of five main components.
-At the core of the system is WallFacer Attention, which leverages multiple ring-style P2P (peer-to-peer) communication strategies to enhance the efficiency of distributed attention computation.
-The Communication Configuration Generator plays a critical role in initially assigning keys and values to their corresponding ranks.
-Meanwhile, the Communication Topology Scheduler outlines the placement of parallelism across various computing nodes.
-The Dataloader is designed to organize tokens within each sub-sequence according to their mask (causal or full) and distribute them across different GPUs.
-Finally, the WallFacer Runtime incorporates additional supporting techniques for the training process, such as gradient checkpointing.
-We will now explore the details of these components in depth.
-3.2 WallFacer Attention
-Figure 4: An example of WallFacer Attention on 64 GPUs.
-Each team member forms a sub-ring with members with the same local rank from other teams in the same ring communication group, reducing the communication volume of each team member by 75%.
+To solve these drawbacks, we develop the StarTrail training system, which we will detail in the following section.
+3.2 StarTrail Attention
+Figure 5: Meanings of the symbols that are used in this paper
+$P$ The number of GPUs
+$C$ The parallel size of StarTrail (team size)
+$H$ The hidden dimension size of the Transformer blocks
+NN The total number of tokens within the whole sequence
+$B$ The training batch size
+$W$ The communication bandwidth between GPUs
+$L$ The communication latency between GPUs
 As discussed in the previous section, a major limitation of Ring Attention is the extensive amount of peer-to-peer (P2P) communication required, which becomes problematic in environments with weak connections between computing nodes.
 We enhance the ring sequence parallelism by introducing an additional dimension.
-This is achieved by duplicating the queries, keys, and values, thus dividing the communication tasks and distributing them within each team.
-This segmentation markedly improves computational efficiency and scalability.
-We will now go through the specifics of the WallFacer Parallelism and provide a theoretical analysis demonstrating its advantages over Ring Attention.
-3.2.1 Training Process of WallFacer Parallelism
-In the WallFacer system, GPUs are grouped into Teams to coordinate computation and communication tasks more efficiently, unlike the isolated operations in traditional settings.
-The queries, keys, and values are gathered within the team, so each member of the team holds the same activation of the whole team before attention.
-WallFacer introduces an additional parameter, $C$, which determines the replication factor of the input and, consequently, the number of GPUs within each team.
+The fundamental idea of StarTrail is akin to a divide-and-conquer strategy.
+During attention, each token must compute its attention score with every other token in the sequence.
+While Ring Attention passes keys and values along a ring of $P$ GPUs over P1P-1 iterations, our approach introduces the concept of a team.
+In this setting, each team member interacts only with a designated portion of the overall sequence, and the results are later aggregated using collective communication.
+Thus, StarTrail can be devided into three phases: preprocessing, ring-phase, and postprocessing.
+For preprocessing, we duplicate the queries within a team using an all-gather operation.
+This ensures that when a GPU receives new keys and values, it can compute the attention scores for the entire team’s queries.
+Similarly, gathering the keys and values allows us to reduce the number of P2P communication iterations by transmitting longer sequences per iteration.
+Following the preprocessing, we enter the ring-style communication phase.
+With the number of GPUs in one team being $C$, $CN/P$ tokens are exchanged in each iteration, and each GPU is responsible for computing $N/C$ tokens.
+This leads to a number of iterations of $\frac{N/C}{CN/P}=\frac{P}{C^{2}},$ within a smaller ring, which we refer to as a subring.
+For convenience, we group $\frac{P}{C^{2}}$ adjacent teams into a team group for subring communication, where GPUs sharing the same local team rank form the subring.
+An initial P2P communication step is executed to ensure that each team group has access to the complete set of keys and values for the sequence (details are provided in the Appendix).
+After completing the subring iterations, each GPU holds $1/C$ of the overall computation result for its team.
+With the help of online softmax, we then apply a simple reduce-scatter to combine these results while eliminating the duplicate tokens, which we refer to as the postprocessing.
+Throughout the attention process, asynchronous communication is employed alongside the early launch of communication kernels to maximize the overlap of computation and communication tasks.
+Now we will delve into more details in the StarTrail training process.
+3.2.1 Configurations of StarTrail Parallelism
+In the StarTrail system, GPUs are grouped into Teams to coordinate computation and communication tasks more efficiently.
+StarTrail introduces an additional parameter, $C$, which determines the replication factor of the input and, consequently, the number of GPUs within each team.
 The range of $C$ is from 1 to $\sqrt{P}$.
@@
 When $1<C<\sqrt{P}$, it becomes a structure with multiple rings looping concurrently.
-There are three main distinctions between the operational processes of WallFacer and Ring Attention: QKV_matmul & all-gather, attention iteration, and reduce-scatter.
-Throughout the attention process, asynchronous communication is employed alongside the early launch of communication kernels to maximize the overlap of computation and communication tasks.
-The complete forward process of a WallFacer Transformer block is outlined in Algorithm 1.
-Algorithm 1 WallFacer Attention Block (Forward)
+Algorithm 1 StarTrail Attention Block (Forward)
 0: Input sequence x, Linear Function query, key, and value, attention parallelism size c, global rank r, global size gs, team process group pg, init send/recv target $\textbf{r}_{\mathrm{send}}$ and $\textbf{r}_{\mathrm{recv}}$
@@
 3: get the ring P2P target $\textbf{r}_{\mathrm{next}}$ and $\textbf{r}_{\mathrm{last}}$ with get_P2P_ranks(r, gs, c)
-4:  initialize attention score O, extra statistics lse to zero.
-5:  for $1\leq i\leq$world_size$/c^{2}$ do
+4: initialize attention score O, extra statistics lse to zero. // lse stands for log-sum-exp
+[…]
 6:  wait for $\textbf{req}_{\mathrm{send}}$ and $\textbf{req}_{\mathrm{recv}}$
@@
 Forward Propagation.
-In Figure 4, we have an example of one team of four GPUs out of all the 64 GPUs performing WallFacer-style attention.
+In Figure 4, we have an example of one team of four GPUs out of all the 64 GPUs performing StarTrail-style attention.
 Each training iteration begins with the dataloader splitting the entire input sequence of length $N$ into $N/P$ sub-sequences, which are then loaded onto each GPU.
@@
 These are computed separately via matrix multiplication, followed immediately by the launch of the all-gather kernel, which gathers the above QKVs within the team, allowing for the overlap of up to two-thirds of the communication with computation.
+Figure 6: An example of ring initialization process of 8GPUs and 4 sub-rings in StarTrail.
 Once this phase is complete, each GPU within the team possesses the same Q, K, and V, each of a length of $\frac{CN}{P}$.
-To distribute the communication and computation tasks among the team members, we divide the original workload based on four specific ranks assigned to each GPU by the Communication Configuration Generator.
-These ranks determine each GPU’s partners and position within the P2P ring, details of which will be discussed in Section 3.3.
-Following the setup, the Keys and Values are dispatched to their designated locations within the cluster to establish the initial sub-ring, setting the stage for the multi-ring iteration phase of WallFacer attention.
-Given that each sub-sequence is $\frac{CN}{P}$ long and each GPU is tasked with computing the attention score for $\frac{1}{C}$ of the whole sequence, it results in $\frac{N/C}{CN/P}=P/C^{2}$ rounds of communication.
+To distribute the communication and computation tasks among the team members, we divide the original workload based on four specific ranks assigned to each GPU.
+These ranks determine each GPU’s partners and position within the P2P ring, as is shown in Figure 6.
+Following the setup, the Keys and Values are dispatched to their designated locations within the cluster to establish the initial sub-ring, setting the stage for the multi-ring iteration phase of StarTrail attention.
+Given that each sub-sequence is $\frac{CN}{P}$ long and each GPU is tasked with computing the attention score for $\frac{1}{C}$ of the whole sequence, it results in $\frac{N/C}{CN/P}-1=P/C^{2}-1$ rounds of communication.
 This implies that there are $P/C^{2}$ GPUs in one ring.
-The iteration process, conducted in a Flash-Attention style, involves storing the log-sum-exp (lse) and intermediate output O, which are updated step by step.
+The iteration process involves storing the log-sum-exp (lse) and intermediate output O, which are updated step by step.
 Queries are retained locally, while Keys and Values circulate through the ring via P2P communication.
@@
 Each GPU ultimately contains the final attention score for its portion of the sequence over the entire sequence.
-The output of this forward attention block is finalized after a standard LayerNorm and FeedForward layer process.
 Backward Propagation.
@@
 The inner loop, however, handles the gradients for queries, which start initialized as zero and are circulated along the sub-rings together with the Queries themselves.
-During each iteration, the approach mirrors the backward computation method used in flash-attention[dao2022flashattention], where the updated gradient of the current query shard is passed to the next GPU in the ring, while the gradients for keys and Values are retained for subsequent query shards.
-Initial communication ranks for sending and receiving are still configured by the Communication Configuration Generator.
-Additionally, there is an extra P2P communication required at the end of the process to send the query and its gradient back to their original location following the loop.
+During each iteration, the approach mirrors the backward computation method used in FlashAttentiondao2022flashattention , where the updated gradient of the current query shard is passed to the next GPU in the ring, while the gradients for keys and Values are retained for subsequent query shards.
 3.2.2 Theoretical Analysis
-In this section, we will discuss two major questions: how much communication reduction WallFacer can bring compared with Ring Attention, and how much extra memory cost it comes with.
-During the analysis, we will employ a case study using the WallFacer system with an attention parallel size of $C$ = 4 on a llama-30B model, which consists of 64 layers.
-For this model, referred to as model M, the batch size = $B$ is set to 1, the sequence length = $N$ to 65, 536, the hidden dimension = $H$ to 6, 656, and the number of GPUs = $P$ to 64.
+During the analysis, we will employ a case study using the StarTrail system with an attention parallel size of $C$ = 4 on a llama-30B model, which consists of 64 layers.
+For this model, referred to as model M, the batch size = $B$ is set to 1, the sequence length = $N$ to 65536, the hidden dimension = $H$ to 6656, and the number of GPUs = $P$ to 64.
 Additionally, the computation will utilize bfloat16 precision.
@@
 For Ring Attention, the communication is primarily due to the ring P2P loop.
-As the total number of iterations done is $P$, the total communication overhead can be calculated as:
-$p(\frac{2BNH}{PW}+L)=\frac{2BNH}{W}+PL$ (2)
+As the total number of iterations done is P1P-1, the total communication overhead can be calculated as:
+$(P-1)(\frac{2BNH}{PW}+L)=\frac{2BNH(P-1)}{WP}+(P-1)L$ (1)
 and this overhead can be partially overlapped with the attention computation.
-For WallFacer, the communication overhead comes from both collective and P2P. The collective overhead for all-gather and reduce-scatter is:
+For StarTrail, the communication overhead comes from both collective and P2P. The collective overhead for all-gather and reduce-scatter is:
 $\frac{4BNH(C-1)}{PW}$ (2)
 while the P2P communication can be similarly computed as:
-$\frac{P}{C^{2}}(\frac{2CBNH}{PW}+L)=\frac{2BNH}{CW}+\frac{P}{C^{2}}L$ (4)
-The advantages of WallFacer over Ring Attention during the ring-P2P phase are evident in three main aspects: 1) Reduced Communication and Latency: Ring Attention requires C times more communication than WallFacer, significantly increasing the bandwidth requirement across the entire cluster.
-For the llama 30B model M, the total communication volume of ring P2P communication and collective communication volume for Ring Attention and WallFacer can be computed as 1.625 GB and 0.152 GB(collective) + 0.406GB (P2P) = 0.558GB.
-Furthermore, while Ring Attention necessitates $P$ iterations per attention block, WallFacer only requires $\frac{P}{C^{2}}$, reducing the latency overhead by $C^{2}$.
-2) Localized Communication: In scenarios like those depicted in Figure 3, WallFacer’s ring P2P communication can be confined within the same computing node, where bandwidth is typically much higher than between computing nodes.
+$(\frac{P}{C^{2}}-1)(\frac{2CBNH}{PW}+L)=\frac{(P-C^{2})2BNH}{CPW}+(\frac{P}{C^{2}}-1)L$ (3)
+The advantages of StarTrail over Ring Attention during the ring-P2P phase are evident in three main aspects: 1) Reduced Communication and Latency: Ring Attention requires C times more communication than StarTrail, significantly increasing the bandwidth requirement across the entire cluster.
+For the llama 30B model M, the total communication volume of ring P2P communication and collective communication volume for Ring Attention and StarTrail can be computed as 1.625 GB and 0.152 GB(collective) + 0.406GB (P2P) = 0.558GB.
+Furthermore, while Ring Attention necessitates P1P-1 iterations per attention block, StarTrail only requires $\frac{P}{C^{2}}-1$, reducing the latency overhead by around $C^{2}$.
+2) Localized Communication: In scenarios like those depicted in Figure 3, StarTrail’s ring P2P communication can be confined within the same computing node, where bandwidth is typically much higher than between computing nodes.
 Conversely, Ring Attention demands inter-node communication during every iteration, which can be less efficient.
-3) Enhanced Overlap of Communication and Computation: During each iteration, the communication volume of WallFacer is $C$ times higher than that of Ring Attention, while the computational volume during attention is approximately $C^{2}$ times greater.
-This higher computation-to-communication ratio makes it easier for WallFacer to overlap P2P communication with computation, enhancing overall efficiency.
-Additionally, the collective communication in WallFacer is minimal due to: 1) Efficient Overlapping: The all-gather communication overlaps significantly with the QKV matrix multiplication, and the reduce-scatter communication partially coincides with the final attention score update.
-2) Scale Considerations: Compared to the P2P communication column, there is a P in the denominator, which is substantial during large-scale training.
-This implies that collective communication constitutes a very small portion of the total communication volume.
+3) Enhanced Overlap of Communication and Computation: During each iteration, the communication volume of StarTrail is $C$ times higher than that of Ring Attention, while the computational volume during attention is approximately $C^{2}$ times greater.
+This higher computation-to-communication ratio makes it easier for StarTrail to overlap P2P communication with computation, enhancing overall efficiency.
 Memory Analysis.
 In this section, we estimate the theoretical peak memory requirements necessary to store the model weights, activations, and optimizer states.
-Our implementation utilizes the Adam Optimizer [kingma2017adam], bfloat16 precision, and Zero-2 optimization [rajbhandari2020zero].
-Since the WallFacer architecture does not alter the model weights or the optimizer, we can assume that the memory costs associated with the model and the optimizer remain constant across both methods.
+Our implementation utilizes the Adam Optimizer kingma2017adam , bfloat16 precision, and Zero-2 optimization rajbhandari2020zero .
 We name the memory cost for the model and optimizer as $M_{m+o}$.
@@
 $A=\frac{B\times N\times H}{P}$ (4)
-As we use the checkpointing scheme from [li2024distflashattn], a model of $Y$ layers needs to save $Y+1$ activations as checkpoints.
+As we use the checkpointing scheme from li2024distflashattn , a model of $Y$ layers needs to save $Y+1$ activations as checkpoints.
 Now we calculate the approximate peak memory after Q, K, and V are already calculated and before the attention computation at the last layer of the whole model.
-For Ring Attention and WallFacer, the peak memories are:
-$PM_{Ring}=M_{m+o}+(Y+1)A+3A=M_{m+o}+(Y+4)A$ (6)
-$PM_{Wall}=M_{m+o}+(Y+1)A+3CA=M_{m+o}+(Y+3C+1)A$ (7)
-, where C is the WallFacer attention dimension.
-And for the example model M, the peak memory would be $M_{m+o}+68A$ and $M_{m+o}+77A$, and the extra memory cost compared with Ring Attention is less than 13.2%, while the P2P communication volume is reduced by about 75%.
+For Ring Attention and StarTrail, the peak memories are:
+$PM_{Ring}=M_{m+o}+(Y+4)A$ (5)
+$PM_{Star}=M_{m+o}+(Y+3C+1)A$ (6)
+, where C is the StarTrail attention dimension.
+And for the example model M, the peak memory would be $M_{m+o}+68A$ and $M_{m+o}+77A$, and the extra memory cost compared with Ring Attention is a lot less than 13.2%, while the P2P communication volume is reduced by about 75%.
 In a word, the extra memory cost is acceptable as a tradeoff for the communication reduction.
-3.3 Communication Configuration Generator
-The Communication Configuration Generator, as introduced in the previous section, plays a crucial role in the initial setup of queries, keys, and values on GPUs and determining each GPU’s position within the sub-rings.
-Figure 5: An example of ring initialization process of 8GPUs and 4 sub-rings in the Communication Configuration Generator
+4 Evaluation
+Table 1: Cluster and Model Configurations. All GPUs are connected with NVLink with computing nodes.
+GPU dev. $\times$ node Mem. (GB) inter-node bandwidth
+H100 8 $\times$ 8 80 8*400Gbps InfiniBand
+A100 16 $\times$ 2 40 100Gbps Ethernet
+A100 8 $\times$ 4 40 100Gbps Ethernet
+A100 4 $\times$ 8 40 100Gbps Ethernet
+Model #Heads #Layers Dim.
+GPT 3B 12 16 4096
+GPT 7B 32 32 4096
+DiT 1B 24 24 1536
+The computational resources we use in the experiments include a local Nvidia H100 cluster with eight nodes and three Nvidia A100 clusters, as listed in table 1.
+We utilize two model types of total three settings, as listed in table 1.
+For the DiT(Diffusion Transformer) model, we use similar configurations as those in Stable Diffusion 3esser2024sd3 .
+We utilize the backbone Diffusion Transformer only, without other components like the text and image encoders.
+During training, both models use bfloat16 precision and a batch size of 1 to accommodate longer input sequences.
+In the evaluation section, we aim to answer three major questions: 1) How much improvement in throughput can StarTrail bring?
+Additionally, how adaptable is StarTrail to clusters with both good and poor inter-node connections?
+2) Is the additional memory cost incurred by StarTrail acceptable considering the throughput improvement it offers?
+3) How does StarTrail perform in scenarios of weak and strong scaling?
+Specifically, does it outperform Ring Attention when scaled to handle longer inputs?
+Figure 7: Throughput evaluation of Ring Attention and StarTrail on 32 GPUs from three different clusters.
+We place the performance of StarTrail with both C=2 and C=4 in the figure.
+The configurations are marked in the titles of the sub-figures.
+For instance, A100_40GB_8(1B, 512K) represents that the experiment is on machines with 8 Nvidia A100 40GB GPUs in each node, the model used has one billion parameters, and the sequence length is 512k.
+4.1 Throughput and Adaptability
+Our first experiment aims to assess the performance of StarTrail and Ring Attention across different clusters with varying environments, testing the adaptability of both methods.
+There are several factors influencing the efficiency of ring-style attention computation:
+Theoretical Computation-Communication Volume Ratio: Primarily determined by the sequence length used during training.
+Attention computation exhibits a computational complexity of $O(N^{2}\cdot H)$, whereas P2P communication complexity is $O(N\cdot H)$.
+Thus, the model configuration does not impact this ratio; only the sequence length does.
+A larger $N$ increases the computation-communication ratio, facilitating easier overlap of communication with computation.
+Compute Capability and Connectivity of GPUs: The computing overhead, given a specific volume, affects the computation-communication overhead ratio.
+Higher compute capabilities make overlapping more challenging.
+We utilize two sets of GPUs in this evaluation: Nvidia A100 40GB and Nvidia H100 80GB, with the latter offering significantly higher theoretical tflops on bf16 computations.
+Connectivity is considered in two parts: intra-node and inter-node.
+Our clusters are equipped with NVLink, ensuring robust intra-node communication.
+For inter-node communication, our H100 nodes leverage InfiniBand with eight adapters per node for superior inter-node bandwidth, whereas the Google Cloud servers use Ethernet.
+The diversity in node configurations (8-GPU and 16-GPU nodes) allows us to assess adaptability across different topologies.
+This evaluation not only highlights the inherent differences between the schemes but also tests their flexibility in various hardware settings.
+The results of our evaluation are illustrated in Figure 7.
+We measure throughput in thousands of tokens per second.
+To better demonstrate how to select the optimal configuration of StarTrail under each condition, we included two configurations, Star-2 and Star-4, in the figure.
+We omit configurations with lower performance for clarity.
+As indicated in the figure, in all six settings, at least one configuration of StarTrail achieves higher throughput than Ring Attention, with 2.114x, 1.414x, 1.629x, 1.425x, 1.360x, 1.199x, 1.771x, and 1.346x the throughput of Ring Attention.
+This advantage is primarily due to the additional parallel dimension that StarTrail introduces.
+Unlike Ring Attention, which requires inter-node P2P communication in each iteration, StarTrail’s P2P communication is mostly confined intra-node, except for initial data transfers.
+This experiment clearly demonstrates StarTrail’s superior performance across various environments.
+Another observation is that the optimal configuration for StarTrail may vary depending on the environment, reflecting differences in the computation-communication ratio and the trade-offs between collective and P2P communication.
+Figure 8: The normalized relative memory cost of different configurations of StarTrail compared with Ring Attention on different clusters.
+4.2 Memory Consumption
+The memory results, displayed in Figure 8, reveal that for the configurations yielding the highest throughput, StarTrail consumes between 7.9% and 30.79% more GPU memory than RingAttention, while achieving 1.199x to 2.114x throughput.
+Moreover, in scenarios involving larger models, the relative increase in memory consumption due to QKV duplication diminishes.
+This reduction is explained by equation 6 This phenomenon is further evidenced by the fact that the extra memory ratio for experiments with the 7B model is significantly smaller than that for the 3B and 1B model.
+Considering the substantial throughput gains provided by StarTrail, this tradeoff between memory usage and efficiency is deemed acceptable.
+(a) Strong scaling experiments with fixed sequence length of 128K.
+4.3 Strong and Weak Scaling
+In the scaling tests we carry out experiments for both strong and weak scaling.
+For strong scaling, we fix the sequence length to 128K while increasing the number of GPUs from 8 to 64 for the GPT model and from 8 to 32 for the DiT model.
+For weak scaling, we scale the sequence length from 128k to 512k for the DiT model and from 64k to 512k for the GPT model proportionally increasing the number of GPUs from 8 to 32.
+As is depicted in Figure 10(a) and 10(b), StarTrail shows obvious advantage over Ring Attention as we increase the number of GPUs.
+The results for strong scaling can also explained by the computation-communication ratio.
+When scaled to more GPUs, the local sequence length on each GPU becomes smaller, and as explained the previous sections, makes it harder to overlap the P2P communication with attention computation.
+The overall scaling performance is limited by the nature of ring-style communication, but we still consider our improvement over Ring Attention meaningful due to the necessity of using Ring-style Parallelisms during training.
+(b) Weak scaling Experiments
+In summary, StarTrail shows better scalability in both strong and weak scaling experiments, making it a better choice for large-scale Transformer model training.
+5 Conclusion
+StarTrail represents an advanced near-infinite-context Transformer model training system, featuring a communication-optimized concentric ring sequence parallelism scheme.
+Through experiments, we demonstrate that our system not only achieves high efficiency across various training environments but also excels under both strong and weak scaling conditions for both CV and NLP models.
+Current limitations of StarTrail include that although orthogonal, we can still further improve the co-design of StarTrail and hybrid parallelism in future works.
+In an era increasingly demanding longer contexts for both NLP and CV, StarTrail is poised to make significant contributions to the industry and inspire innovative research in academia.
+6 acknowledgement
+Yang You’s research group is being sponsored by NUS startup grant (Presidential Young Professorship), Singapore MOE Tier-1 grant, ByteDance grant, NUS ARTIC grant, Apple grant, Alibaba grant, Google Research and Google grant for TPU usage.
+NeurIPS Paper Checklist
+The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact.
+Do not remove the checklist: The papers not including the checklist will be desk rejected.
+The checklist should follow the references and follow the (optional) supplemental material.
+The checklist does NOT count towards the page limit.
+Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
+  • You should answer [Yes] , [No] , or [N/A] .
+  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
+  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
+The checklist answers are an integral part of your paper submission.
+They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers.
+You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
+The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation.
+While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used").
+In general, answering "[No] " or "[N/A] " is not grounds for rejection.
+While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate.
+All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix.
+If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
+IMPORTANT, please:
+  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
+  • Keep the checklist subsection headings, questions/answers and guidelines below.
+  • Do not modify the questions and only use the provided macros for your answers.
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: We have clearly claimed that we improved the efficiency of ring-style sequence parallelism and provide analysis and experiments in the following sections.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: As mentioned in the conclusion, although orthogonal to other parallelism, we have not consider the co-design with other parallelism for better performances.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: We do not have any theoretical result.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: We have provided the open-source models and GPU configurations we used in the experiments. The algorithm of the parallelism is also provided in this paper.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [No]
+  24. Justification: Our code is not currently available publicly, but we will prepare for open-source if accepted.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: We have provided all the details in the experiment sections.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [No]
+  34. Justification: Our experiment results are measured by measuring a large number of training iterations and taking the average. It is not suitable for our case to prove statistical significance.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: All details are mentioned in the experiment section.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We reviewed the NeurIPS Code of Ethics and this research conforms with the guidelines.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [N/A]
+  49. Justification: This work is simply improving the training efficiency and has no societal impact.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: We do not introduce any new model or dataset.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We have credited the assets properly.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: Our work does not release new assets.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: Our research does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: Our research does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: Our key method is modifying the parallelism technique and is not related with LLM itself.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+Appendix A Technical Appendices and Supplementary Material
+Figure 10: A comparison between naive and zigzag dataloader for 8 GPUs with attention parallel dimension of 2.
+The corresponding initialization can be found in Figure 6 with the same configuration.
+The improvement of efficiency from load-balancing increases with the number of GPUs.
+A.1 Sequence Parallelism Dataloader
+The causal masks present a challenge due to the unbalanced computational load across GPUs: sub-sequences at the beginning of a sequence require significantly more computation than those at the end.
+To address this imbalance and achieve load equilibrium among GPUs, we modify the ZigZag scheme introduced by [42], illustrated in Figure 10.
+The figure illustrates the simplest case of zigzag load-balancing.
+Notably, the effectiveness of this strategy improves as the number of GPUs increases.
+This improvement correlates with the expanding difference in computation volume between the first and the last token, which escalates as the sequence length extends.
+This approach ensures that the total workload on each GPU is balanced, eliminating the need for additional communication mechanisms like those employed in DistFlashAttention[20].
+A.2 Details in the training process
+Apart from the attention process described in Algorithm 1, we would like do provide a few other details to be more comprehensive.
+First, before the concentric ring attention, we need to initialize the the Keys and Values and determine each GPU’s position within the rings.
 Initially, for the setup stage, it is essential to establish the sub-rings by rearranging the activation positions.
 Specifically, during forward propagation, the queries do not require rearrangement; however, the keys and values must be transmitted to their corresponding positions in the ring prior to commencing the loop.
-As illustrated in Figure 5, this initialization ensures that each team member holds a different shard of keys and values.
+As illustrated in Figure 6 and algorithn, this initialization ensures that each team member holds a different shard of keys and values.
 Moreover, it guarantees that no two teams within the same ring possess identical keys and values.
-Furthermore, Team 0 and Team 1 form what is termed a team group, where ring P2P communication occurs exclusively within this group, similarly for Team 2 and Team 3.
-A team group is defined as a collection of teams that participate in the same sub-rings.
-The number of teams within a team group can be calculated as $Team\ size=\frac{P}{C^{2}}$.
-We give the details of the determination of initial sending targets in the case of placing teams within the same computing nodes in Algorithm 2, and the receiving source can also be calculated similarly.
+Initially, for the setup stage, it is essential to establish the sub-rings by rearranging the activation positions.
+Specifically, during forward propagation, the queries do not require rearrangement; however, the keys and values must be transmitted to their corresponding positions in the ring prior to commencing the loop.
+As illustrated in Figure 6 and algorithm, this initialization ensures that each team member holds a different shard of keys and values.
+Moreover, it guarantees that no two teams within the same ring possess identical keys and values.
 Algorithm 2 get_init_send()
@@
 3: target team = target team group rank * team group size + $\textbf{r}_{t}$ // $\textbf{d}_{a}$
-4:  target device intra-team rank = $\textbf{r}_{t}$ $\textbf{d}_{a}$
+4: target device intra-team rank = $\textbf{r}_{t}$ % $\textbf{d}_{a}$
 5: target global rank = target team * $\textbf{d}_{a}$ + target device intra-team rank
 6: return target global rank
-After the initialization of activations, we can set up the rings by providing the GPUs their last and next GPU within their rings, as is described in Algorithm 3
+After the initialization of activations, we can set up the rings by providing the GPUs their last and next GPU within their rings, as is described in Algorithm 3.
 Algorithm 3 get_P2P_config()
@@
 2: self team group rank = $\frac{r_{t}}{team\ group\ size}$
-3:  next team in group = $(r_{t}+1)\%$ team group size $+$ team group size $\times$×self team group rank
-4:  last team in group = $(r_{t}-1)\%$ team group size $+$ team group size $\times$×self team group rank
-5:  next device global rank = $r_{a}+$ next team in group ×$\times d_{a}$
-6:  last device global rank = $r_{a}+$ last team in group ×$\times d_{a}$
+3: next team in group = $(r_{t}+1)\%$ team group $size ×\timesself team group$ rank
+4: last team in group = $(r_{t}-1)\%$ team group $size ×\timesself team group$ rank
+5: next device global rank = $r_{a}+$ next team in group $\times d_{a}$
+6: last device global rank = $r_{a}+$ last team in group $\times d_{a}$
 7: return next device global rank, last device global rank
-3.4 Communication Topology Scheduler
-The communication topology scheduler employs a grid-search algorithm to discover the optimal parallelism configuration, tailored to the specifics of the current cluster, model, and input data.
-When using Ring Attention, all GPUs form a single ring, which considerably limits the flexibility of adjusting the placement scheme.
-In contrast, WallFacer expands the tuning space with its unique multi-ring structure in two significant ways.
-First, the ring configuration in WallFacer is determined by the parallelism dimension $C\in[1,\sqrt{P}]$.
-Smaller $C$ results in less collective communication, a higher total communication volume, and more inter-node communication, while higher C leads to more communication volume during all-gather and reduce-scatter, which may exceed the overhead of the QKV matrix multiplication that we use for overlapping, but can highly reduce the overall communication volume.
-Second, the additional dimension provided by WallFacer allows for two strategies in parallelism placement.
-One strategy is to keep the all-gather and reduce-scatter operations intra-node, which confines these processes within the same computing node.
-The other strategy is to keep the ring P2P communication intra-node, reducing latency and potentially enhancing performance.
-The choice between these strategies depends on the extent of overlapping achievable within the ring and the communication volume during collective communications.
-Based on the above tuning space, we provide an auto-scheduler, grid-searching through all possible configurations for the one that provides the highest throughput, which can be described as:
-$\begin{split}Config=\arg\max_{C,placement}(Profile(C\in[1,\sqrt{P}],\\ placement\in[P2P\_intra,Collect\_intra])\end{split}$ (8)
-The scheduler requires only a few iterations to profile the performance of automatically generated configurations.
-The time spent on this profiling is negligible compared to the entire training process, as it only needs to be performed once for each new computing cluster used for training.
-Figure 6: A comparison between naive and zigzag dataloader for 8 GPUs with attention parallel dimension of 2.
-The corresponding initialization can be found in Figure 5 with the same configuration.
-The improvement of efficiency from load-balancing increases with the number of GPUs.
-3.5 Sequence Parallelism Dataloader
-To accommodate both full mask and causal mask configurations for self-attention, we have implemented two distinct data loading schemes for load-balancing.
-For full masks, sequences are straightforwardly divided into sub-sequences of equal lengths and distributed among the devices.
-However, causal masks present a challenge due to the unbalanced computational load across GPUs; sub-sequences at the beginning of a sequence require significantly more computation than those at the end.
-To address this imbalance and achieve load equilibrium among GPUs, we modify the ZigZag scheme introduced by [zilin2024ringflashattention], illustrated in Figure 6.
-The figure illustrates the simplest case of zigzag load-balancing.
-Notably, the effectiveness of this strategy improves as the number of GPUs increases.
-This improvement correlates with the expanding difference in computation volume between the first and the last token, which escalates as the sequence length extends.
-This approach ensures that the total workload on each GPU is balanced, eliminating the need for additional communication mechanisms like those employed in DistFlashAttention[li2024distflashattn].
-Figure 7: Throughput evaluation of Ring Attention and WallFacer on 32 GPUs from three different clusters.
-We place the performance of WallFacer with both C=2 and C=4 in the figure.
-The configurations are marked in the titles of the sub-figures.
-For instance, A100_40GB_8(1B, 512K) represents that the experiment is on machines with 8 Nvidia A100 40GB GPUs in each node, the model used has one billion parameters, and the sequence length is 512k.
-3.6 WallFacer Runtime
-WallFacer is written in PyTorch[pytorch] and uses the PyTorch torch.autograd.function and NCCL[nccl] backend for forward and backward implementation.
-WallFacer also employs multiple techniques during runtime to improve its overall training efficiency.
+A.3 StarTrail Runtime
+StarTrail is written in PyTorch[29] and uses the PyTorch torch.autograd.function and NCCL[28] backend for forward and backward implementation.
+StarTrail also employs multiple techniques during runtime to improve its overall training efficiency.
 Ingetrate Flash Attention.
-The WallFacer attention mechanism involves multiple iterations that loop over Keys and Values, with each iteration still using traditional self-attention with corresponding Query, Key, and Value (QKV).
-This approach enables WallFacer to incorporate flash attention effectively, extending its capability by preserving intermediate states across iterations.
-Additionally, WallFacer enhances the efficiency of the forward process with the help of torch JIT to fuse kernels aside from flash attention.
+The StarTrail attention mechanism involves multiple iterations that loop over Keys and Values, with each iteration still using traditional self-attention with corresponding Query, Key, and Value (QKV).
+This approach enables StarTrail to incorporate flash attention effectively, extending its capability by preserving intermediate states across iterations.
+Additionally, StarTrail enhances the efficiency of the forward process with the help of torch JIT to fuse kernels aside from flash attention.
 Overlap communication with computing.
-In WallFacer attention, P2P communication and self-attention computing are interleaved across iterations, each incurring considerable time.
-To mitigate this, WallFacer employs a double buffering technique to asynchronously execute communication and computing kernels, effectively overlapping these processes and enhancing GPU utilization.
+In StarTrail attention, P2P communication and self-attention computing are interleaved across iterations, each incurring considerable time.
+To mitigate this, StarTrail employs a double buffering technique to asynchronously execute communication and computing kernels, effectively overlapping these processes and enhancing GPU utilization.
 Save recomputation with checkpoints.
-WallFacer adopts the checkpointing strategy introduced by DistFlashAttn[li2024distflashattn], placing checkpoints at the end of the self-attention phase rather than the FFN of each transformer layer.
+StarTrail adopts the checkpointing strategy introduced by DistFlashAttn[20], placing checkpoints at the end of the self-attention phase rather than the FFN of each transformer layer.
 This checkpoint placement effectively obviates the need to recompute the self-attention forward process during the backward pass, avoiding redundant attention computation.
-4 Evaluation
-The computational resources we use in the experiments include a local Nvidia H100 cluster with eight nodes and three Nvidia A100 clusters, as listed in table 2.
-We utilize two model types of total three settings, as listed in table 3.
-For the DiT(Diffusion Transformer) model, we use similar configurations as those in Stable Diffusion 3[esser2024sd3].
-We utilize the backbone Diffusion Transformer only, without other components like the text and image encoders.
-During training, both models use bfloat16 precision and a batch size of 1 to accommodate longer input sequences.
-Table 2: Cluster Configurations. All GPUs are connected with NVLink with computing nodes. #GPU represents local size times the number of nodes.
-GPU #GPUs Memory Inter-node
-Nvidia H100 8*8 80GB InfiniBand
-Nvidia A100 16*2 40GB 100Gbps Ethernet
-Nvidia A100 8*4 40GB 100Gbps Ethernet
-Nvidia A100 16*2 40GB 100Gbps Ethernet
-Table 3: Model Configurations
-Model Name #Heads #Layers Hidden Dim
-GPT 3B 12 16 4096
-GPT 7B 32 32 4096
-DiT 1B 24 24 1536
-In the evaluation section, we aim to answer three major questions:
-  • How much improvement in throughput can WallFacer bring? Additionally, how adaptable is WallFacer to clusters with both good and poor inter-node connections?
-  • Is the additional memory cost incurred by WallFacer acceptable considering the throughput improvement it offers?
-  • How does WallFacer perform in scenarios of weak and strong scaling? Specifically, does it outperform Ring Attention when scaled to handle longer inputs?
-Figure 8: The normalized relative memory cost of different configurations of WallFacer compared with Ring Attention on different clusters.
-4.1 Throughput and Adaptability
-Our first experiment aims to assess the performance of WallFacer and Ring Attention across different clusters with varying environments, testing the adaptability of both methods.
-There are several factors influencing the efficiency of ring-style attention computation:
-Theoretical Computation-Communication Volume Ratio: Primarily determined by the sequence length used during training.
-Attention computation exhibits a computational complexity of $O(N^{2}\cdot H)$, whereas P2P communication complexity is $O(N\cdot H)$.
-Thus, the model configuration does not impact this ratio; only the sequence length does.
-A larger $N$ increases the computation-communication ratio, facilitating easier overlap of communication with computation.
-We evaluate varying sequence lengths to explore performance under different ratios.
-Compute Capability and Connectivity of GPUs: The computing overhead, given a specific volume, affects the computation-communication overhead ratio.
-Higher compute capabilities make overlapping more challenging.
-We utilize two sets of GPUs in this evaluation: Nvidia A100 40GB and Nvidia H100 80GB, with the latter offering significantly higher theoretical tflops on bf16 computations.
-Connectivity is considered in two parts: intra-node and inter-node.
-Our clusters are equipped with NVLink, ensuring robust intra-node communication.
-For inter-node communication, while InfiniBand offers high bandwidth at a higher commercial cost, some clusters utilize Ethernet for connectivity.
-Specifically, our H100 nodes leverage InfiniBand with eight adapters per node for superior inter-node bandwidth, whereas the Google Cloud servers use Ethernet.
-The diversity in node configurations (8-GPU and 16-GPU nodes) allows us to assess adaptability across different topologies.
-This evaluation not only highlights the inherent differences between the schemes but also tests their flexibility in various hardware settings.
-The results of our evaluation are illustrated in Figure 7.
-We measure throughput in thousands of tokens per second.
-To better demonstrate how to select the optimal configuration of WallFacer under each condition, we included two configurations, Wall-2 and Wall-4, in the figure.
-We omit configurations with lower performance for clarity.
-As indicated in the figure, in all six settings, at least one configuration of WallFacer achieves higher throughput than Ring Attention, with performance improvements of 114.33%, 41.4%, 62.87%, 42.45%, 35.98%, 19.9%, 77.12%, and 34.62%.
-This advantage is primarily due to the additional parallel dimension that WallFacer introduces.
-Unlike Ring Attention, which requires inter-node P2P communication in each iteration, WallFacer’s P2P communication is mostly confined intra-node, except for initial data transfers.
-This experiment clearly demonstrates WallFacer’s superior performance across various environments.
-Another observation is that the optimal configuration for WallFacer may vary depending on the environment, reflecting differences in the computation-communication ratio and the trade-offs between collective and P2P communication.
-For example, the best $C$ value for A100_16 is 2, while for A100_8, it is 4.
-This variation can be attributed to the topological differences: the 16-GPU-node cluster benefits less from reduced P2P communication volume, making a smaller $C$ preferable to decrease collective communication volume while maintaining P2P communication at an acceptable level.
-This finding underscores the importance of the Communication Topology Scheduler we provide, which helps users avoid the complex process of manually analyzing these factors.
-In summary, WallFacer demonstrates superior efficiency and adaptability in all six cases, thanks to its flexible parallelism scheme.
-Table 4: Supported Sequence Length of Ring Attention and WallFacer on one Nvidia A100 80GB GPU.
+Appendix B Additional Experiment
+To comprehensively evaluate the memory consumption of StarTrail and Ring Attention, we compared the maximum supported sequence lengths of StarTrail with those reported in the Ring Attention paper [24].
+As shown in Table 2, although StarTrail requires slightly more memory, it still supports sequence lengths commonly used in training tasks.
+Table 2: Supported Sequence Length of Ring Attention and StarTrail on one Nvidia A100 80GB GPU.
 Supported Seq Len on one 80GB A100 GPU
 (K Tokens)
-Model Size Length Ring Attention WallFacer
+Model Size Length Ring Attention StarTrail
 3B 128
@@
 512
-4.2 Memory Consumption
-Before evaluating the memory consumption of WallFacer, we first compared the supported sequence lengths of WallFacer with those reported in the RingAttention paper [liu2023ring].
-As shown in Table 4, despite the slightly higher memory requirements of WallFacer, it still supports sequence lengths commonly used in training tasks.
-Memory consumption in our experiments stems from both the model weights and activations, with additional costs for WallFacer arising from the duplication of QKV matrices prior to attention computation.
-To quantitatively assess the additional memory required by WallFacer, we monitored the maximum memory allocated by PyTorch [pytorch] during our experiments.
-It is important to note that memory fragmentation in PyTorch can impact memory allocation efficiency.
-To mitigate this, we limited the sequence lengths in our training to prevent PyTorch from triggering cuda_free when the allocated memory approaches the GPU’s limit, which would otherwise introduce significant overhead.
-The results, displayed in Figure 8, reveal that for the configurations yielding the highest throughput, Ring Attention consumes between 7.9% and 30.79% less GPU memory than WallFacer, except the DiT 512k Wall-4 setting.
-However, considering the substantial throughput gains provided by WallFacer, the additional memory usage is deemed acceptable.
-We observed that the maximum supported sequence lengths were not significantly impacted by this additional memory usage.
-Moreover, in scenarios involving larger models, the relative increase in memory consumption due to QKV duplication diminishes.
-This reduction is explained by equation 7, which shows that as the model size increases—owing to more parameters and additional Transformer layers—the proportion of memory consumed by model weights and activations grows, whereas the absolute extra memory incurred by QKV multiplication remains constant.
-This phenomenon is further evidenced by the fact that the extra memory ratio for experiments with the 7B model is significantly smaller than that for the 3B and 1B model.
-4.3 Strong and Weak Scaling
-In the scaling tests we carry out experiments for both strong and weak scaling.
-Strong Scaling maintains the scale of the problem that we are trying to solve while increasing the computing resource that we use to speed it up.
-So in this experiment, we fix the sequence length to 128K while increasing the number of GPUs from 8 to 64 for the GPT model and from 8 to 32 for the DiT model.
-As id depicted in Figure LABEL:Fig.strong, WallFacer shows gradually more advantage over Ring Attention as we increase the number of GPUs.
-When we use 64 GPUs, WallFacer shows 60.71% higher throughput than Ring Attention for the GPT model and 114.33% for the DiT model on 32 A100 GPUs.
-This can also explained by the computation-communication ratio.
-When scaled to more GPUs, the local sequence length on each GPU becomes smaller, and as explained the previous sections, makes it harder to overlap the P2P communication with attention computation.
-This is exactly the main advantage of WallFacer, reduction of total P2P communication volume.
-So WallFacer can give a promising performance when we are trying to speedup the training of a fixed amount of data with more GPUs.
-In our weak scaling experiments using DiT models, we scale the sequence length from 128k to 512k while proportionally increasing the number of GPUs from 8 to 32.
-Assuming a scaling factor of $k$, both the sequence length and the number of GPUs increase by $k$, resulting in a total computation volume that increases by approximately $k^{2}$.
-Since the computational resources also increase by $k$, the theoretical overall throughput (tokens per second) should remain constant during this weak scaling setting.
-As shown in Figure LABEL:Fig.weak, both WallFacer and Ring Attention exhibit this expected behavior, with throughput remaining relatively constant.
-However, WallFacer consistently achieves higher throughput across different numbers of GPUs.
-This is because, although the local computation-to-communication ratio stays the same in weak scaling, Ring Attention requires a higher proportion of inter-node communication as the number of GPUs increases, whereas WallFacer keeps most communication local.
-In summary, WallFacer shows better scalability in both strong and weak scaling experiments, making it a better choice for large-scale Transformer model training.
-5 Discussion
-5.1 FlashAttention3 and Hopper GPUs
-In addition to the original FlashAttention [dao2022flashattention] used in our experiments, FlashAttention3 [shah2024flashattention3fastaccurateattention] has been introduced, specifically designed for Hopper and newer Nvidia GPUs.
+B.1 Discussion
+B.2 Larger Batch Sizes and Models
+We utilized small batches and model sizes in our experiments because these choices do not affect the underlying improvements in communication efficiency and computation-to-communication ratio that StarTrail provides.
+For larger batch sizes, both communication and computation scale proportionally, leaving the overlapping ability unchanged.
+Similarly, while larger models involve more layers or larger hidden sizes, the key attention computations and corresponding ratios remain unaffected.
+Hence, our conclusions naturally extend to scenarios with larger batches and models.
+B.2.1 FlashAttention3 and Hopper GPUs
+In addition to the original FlashAttention [8] used in our experiments, FlashAttention3 [35] has been introduced, specifically designed for Hopper and newer Nvidia GPUs.
 For FP16 precision, which is utilized in this paper, FlashAttention3 achieves a 1.5-2.0x speedup on Hopper GPUs.
 As discussed in Section 4, reducing attention computation overhead results in more P2P communication not being overlapped, further emphasizing the need to reduce communication volume.
-With the increasing adoption of Hopper GPUs, the significance of the WallFacer system will also grow.
-5.2 WallFacer and Other Parallelisms
-Model Parallelism As is well known, tensor parallelism shards activations by attention heads during attention computation, making it easily combinable with WallFacer with minimal effort.
-However, when combined with tensor parallelism, the need for attention heads can limit the scalability of head-based sequence parallel methods like Ulysses.
-Pipeline parallelism, on the other hand, divides the model across layers without altering the computation patterns within Transformer blocks, making WallFacer orthogonal to it.
-Other Sequence Parallelism WallFacer is orthogonal with other attention-head-sharding-based sequence parallelism approaches, such as DeepSpeed Ulysses [jacobs2023Ulysses].
-While Ulysses distributes attention heads across different devices, WallFacer can independently partition activations along the sequence length dimension.
-In future work, we can explore combining WallFacer with DeepSpeed Ulysses to expand the communication scheduling space, harnessing the scalability of WallFacer alongside the efficiency of Ulysses.
-In summary, WallFacer can be seamlessly integrated with other parallel training techniques, enabling the creation of a hybrid distributed training system.
-6 Related Work
+With the increasing adoption of Hopper GPUs, the significance of the StarTrail system will also grow.
+B.2.2 StarTrail and Other Parallelisms
+Model Parallelism As is well known, tensor parallelism shards activations by attention heads during attention computation, making it easily combinable with StarTrail with minimal effort.
+However, when combined with tensor parallelism, the need for attention heads can limit the scalability of head-based sequence parallel methods like DeepSpeed-Ulysses.
+Pipeline parallelism, on the other hand, divides the model across layers without altering the computation patterns within Transformer blocks, making StarTrail orthogonal to it.
+Other Sequence Parallelism StarTrail is orthogonal with other attention-head-sharding-based sequence parallelism approaches, such as DeepSpeed-Ulysses [14].
+While DeepSpeed-Ulysses distributes attention heads across different devices, StarTrail can independently partition activations along the sequence length dimension.
+In future work, we can explore combining StarTrail with DeepSpeed-Ulysses to expand the communication scheduling space, harnessing the scalability of StarTrail alongside the efficiency of DeepSpeed-Ulysses.
+In summary, StarTrail can be seamlessly integrated with other parallel training techniques, enabling the creation of a hybrid distributed training system.
+B.3 Other Related Works
 Attention Optimization.
@@
 To address these challenges within the GPU, several approaches have been devised to reduce both memory and computational requirements.
-Memory-efficient attention[rabe2022selfattention] introduces a straightforward algorithm that requires only $O(1)$ memory relative to the sequence length, with an extension for self-attention that needs only $O(\log n)$ memory.
+Memory-efficient attention[32] introduces a straightforward algorithm that requires only $O(1)$ memory relative to the sequence length, with an extension for self-attention that needs only $O(\log n)$ memory.
 FlashAttention further minimizes I/O overhead and enhances overall efficiency.
-Additionally, optimization methods specifically tailored for inference, such as PagedAttention[kwon2023efficient], are also being developed to improve the efficiency of attention computations.
+Additionally, optimization methods specifically tailored for inference, such as PagedAttention[19], are also being developed to improve the efficiency of attention computations.
 In this work, we utilize FlashAttention within each iteration to reduce the computation overhead.
 Long-Sequence Training Techniques.
-Sequence Parallelism[Li2021SequencePL] was initially introduced to enhance the efficiency of parallel long-sequence training.
-Ring Attention[liu2023ring] improved communication efficiency through memory-efficient methods[rabe2022selfattention], supporting near-infinite sequence lengths.
-DeepSpeed Ulysses[jacobs2023Ulysses] employs attention head splitting to achieve high efficiency, though it is constrained by the number of heads.
-Megatron Sequence Parallelism focuses on reducing memory costs during Tensor Parallelism, while DistFlashAttention[li2024distflashattn] features a load-balance scheme and a novel gradient checkpoint method.
+Sequence Parallelism[21] was initially introduced to enhance the efficiency of parallel long-sequence training.
+Ring Attention[24] improved communication efficiency through memory-efficient methods[32], supporting near-infinite sequence lengths.
+DeepSpeed-Ulysses[14] employs attention head splitting to achieve high efficiency, though it is constrained by the number of heads.
+Megatron Sequence Parallelism focuses on reducing memory costs during Tensor Parallelism, while DistFlashAttention[20] features a load-balance scheme and a novel gradient checkpoint method.
 Our work builds on these innovations, introducing a system that supports large-scale training with an efficient communication scheme.
 Techniques for Distributed Model Training.
-Distributed model training encompasses two primary areas: 1) Memory Management: Various techniques aim to conserve GPU memory during distributed training, such as mixed precision training[micikevicius2018mixed] and the ZeRO series[rajbhandari2020zero].
+Distributed model training encompasses two primary areas: 1) Memory Management: Various techniques aim to conserve GPU memory during distributed training, such as mixed precision training[26] and the ZeRO series[34].
 In this work, we implement ZeRO-2 to manage optimizer states and gradients efficiently.
-2) Hybrid Parallelism: Frameworks like Megatron[megatron] and Colossal AI[bian2021colossal] integrate multiple forms of parallelism.
-There are various existing Parallelism techniques like Pipeline Parallelism[gpipe, dapple, li2021chimera, Liu2023Hanayo] and Tensor Parallelism[shoeybi2019megatron], which can be combined with WallFacer Parallelism to facilitate large-scale training.
-We are also considering the integration of additional frameworks such as [chang2024centauri] to enhance overlapping capabilities in future implementations.
-7 Conclusion
-WallFacer represents an advanced near-infinite-context Transformer model training system, featuring a communication-optimized multi-ring sequence parallelism scheme.
-Through comprehensive adaptability and scaling experiments, we demonstrate that our system not only achieves high efficiency across various training environments but also excels under both strong and weak scaling conditions for both Computer Vision(CV) and Natral Language Processing(NLP) models.
-In an era increasingly demanding longer contexts for both NLP and CV, WallFacer is poised to make significant contributions to the industry and inspire innovative research in academia.
-[…]
+2) Hybrid Parallelism: Frameworks like Megatron[27] and Colossal AI[4] integrate multiple forms of parallelism.
+There are various existing Parallelism techniques like Pipeline Parallelism[13, 10, 23, 25] and Tensor Parallelism[37], which can be combined with StarTrail Parallelism to facilitate large-scale training.
+We are also considering the integration of additional frameworks such as [6] to enhance overlapping capabilities in future implementations.